### v2

# From the CDC to emerging infectious disease publics: The long-now of polarizing and complex health crises

Abstract.
As the COVID-19 pandemic evolved, the Center for Disease Control and Prevention used Twitter to share updates about the virus and safety guidelines, reaching millions instantly, in what we call the CDC public. We analyze two years of tweets, from, to, and about the CDC using a mixed-methods approach to characterize the nature and credibility of COVID-19 discourse and audience engagement. We found that the CDC is not engaging in two-way communication with the CDC publics and that discussions about COVID-19 reflected societal divisions and political polarization. We introduce a crisis message journey concept showing how the CDC public responds to the changing nature of the crisis (e.g., new variants) using “receipts” of earlier, and at times contradictory, guidelines. We propose design recommendations to support the CDC in tailoring messages to specific users and publics (e.g., users interested in racial equity) and in managing misinformation, especially in reaction to crisis flashpoints.
Twitter, Digital publics, Public health crisis, Crisis communication, Public policy discourse, Temporal crisis distance, affordances, change-point detection
copyright: acmlicensedjournal: TOCHIjournalyear: 2025journalvolume: XXjournalnumber: XXarticle: TOCHIpublicationmonth: 4doi: XX.XXXX/XXXXXXXccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
“But for politics and bad beer, the matter would never have been heard of”
Walter Kempster, Milwaukee Health Commissioner (1841-1918)
The use of technology to disseminate public health messages has historically shaped public health outcomes. In the quote above, the matter Kempster describes is the breakdown of civic life in Milwaukee, WI in 1894 due to a smallpox outbreak that got out of control (Leavitt, 2003)[P.187-188]. While Kempster was a capable physician, he overlooked the importance of communication in explaining health risks to the city population and opening communication channels to receive their feedback. A widely different experience describes the effective control of a smallpox outbreak in 1947 in New York City. In that iteration, government organizations used frequent, clear, and simple messages to keep citizens updated through different channels (e.g., newspapers, radio shows, and even vaccination lapels): “Be safe. Be sure. Get vaccinated.” These clear communication techniques were coupled with collaboration with volunteer organizations (e.g., the Red Cross) which provided more channels to different communities in the city (Leavitt, 2003)[P.189]. The communication of public health messages is still a challenge for organizations today. For example, the recent measles outbreak in the U.S. suggests the continued struggle to manage vaccine disinformation, inequitable health care access, and limits to health literacy. These cases raise important questions for human-computer interaction (HCI) scholars on the role of design in organizations’ use of social media to disseminate public health messages and interact with the public.
To investigate these challenges, this study analyzed the discourse of the Centers for Disease Control and Prevention (CDC) during the COVID-19 Pandemic. Taking a longitudinal approach, we examined how the CDC sent and received communication messages across the first two years of the pandemic on Twitter. [^1] Twitter re-branded to X after the collection of the data used in this study and so, we refer to it as Twitter throughout this paper. This exchange of communication messages on Twitter is an example of the Habermasian public sphere - a space for “the coming together of peoples and the discussion of ideas…often related to governance and Democratic ideals.” (Poor, 2005) These publics collapse around “issues” of discussion (e.g., #COVID) (Dahlberg, 2001), or arenas of discussion (e.g., online neighborhood group (Randazzo et al., 2025) or public health institutional account) (Moe, 2023). Therefore, we refer to the area of discourse we analyze as the CDC public.
Twitter is a virtual “public square” (Jenzen et al., 2021) where different ad hoc “publics” can be identified using hashtags (Bruns and Moe, 2014) which are consumed by interested audiences (Livingstone, 2005). They can also be formed around mentions that are searchable by other users (Kountouri and Kollias, 2023). Retweets, quote-tweets, and responses to Tweets are associated with propagating messages in different ways. Given how different affordances of the platform (Faraj and Azad, [n. d.]; Leonardi and Vaast, [n. d.]) can affect the way frames are propagated through retweets and directed through mentions at other members of the public, we ask:
RQ1a: What are the properties of users whose messages are propagated the most?
Given that earlier work shows how the use of low-credibility information sources (Wang et al., 2023; Hobbs et al., 2024), the use of rich media in the message (Jin et al., 2016; Gupta et al., 2013), and the sentiment portrayed in a message (Valdez et al., 2020) all affect the way messages are propagated in a network, we ask:
RQ1b: How do the properties of a message affect the way they are propagated in the CDC public?
Finally, given that earlier work shows how the properties of the users disseminating the messages, specifically their follower-following ratio (signifying their influence in the network) and their verified status (Harris et al., 2024) affect the way their messages are propagated in publics, we ask:
RQ1c: How are messages from verified users and users with high follower-following ratios in the CDC public different from other users?
Crisis management includes official policies and communication received from government entities and major health organizations. Over recent years, there has been a significant increase in government use of social media as a vehicle for disseminating public health crisis information. In turn, there is an increased expectation among the public for this greater level of communication (Hagen et al., 2018). Using social media for crisis management information can be of great use to the public. For instance, during the 2009 H1N1 influenza outbreak, the Alexandria Virginia Health Department informed citizens of where to receive vaccinations using social media (Merchant et al., 2011).
Crisis communication is the process of providing facts to the public about an unexpected emergency, beyond an organization’s control, that involves the organization and requires an immediate response (Zahry et al., 2023). During health crises, information becomes a ‘necessity, not a luxury’. Public health communication can be compared to political persuasion by framing their messages to push a specific goal (Dennison and Geddes, 2019; Entman, 1993). In turn, audiences have ‘interpretive schemata’ that allow them to understand the message. Messaging can be honed by stressing the salience of specific frames (Oppermann and Viehrig, 2011). These frames can become polarized by different stakeholders in the public (Leavitt, 2003). Discursive polarization is affective (e.g., social identity is polarized by framing discourse as “us vs. them”) and ideologically polarized by stressing the disagreement about the definition of the problem at hand (e.g., we are facing a pandemic) and providing evaluative statements (e.g., #maskssavelives) (Brüggemann and Meyer, 2023). Given the centrality of framing to making sense of a public space, we ask:
RQ2a: What are the major frames of discussion in different publics and how are they polarized?
Boyd defines networked publics as “publics that are restructured by networked technologies; they are simultaneously a space and a collection of people.” (Boyd, 2010) Those societal stakeholders are “actors and organizations viewed as having a vested interest in the decision output.” (Crow et al., 2020)[P.2] Hallahan (Hallahan, 2000) breaks stakeholders into four main groups: (1) active stakeholders: have deep knowledge and involvement in the policy process; (2) aware stakeholders: knowledgeable stakeholders who might not be directly affected by policy change (e.g., policy experts and opinion leaders); (3) aroused stakeholders: while they have a stake in policy, they have low levels of knowledge about it; and (4) inactive stakeholders: lacking both knowledge and involvement (Park and Lee, 2020)[P.642]. Given the importance of identifying different archetypes of stakeholders in a potentially polarized public, we ask:
RQ2b: Who are the main stakeholders in our public? How are they polarized around different issues?
Public health crises are often a time of uncertainty, confusion, and frustration. These heightened emotions amplify community discourse, particularly on social media platforms (Pine et al., 2021). This discourse can be split into two categories: communication pertaining to the risk of infection, and the risks associated with the authorities’ crisis management measures (Gui et al., 2018). As public health crises progress over time, the responses to the public health emergency change, both on the side of public health organizations like the CDC, and from publics (Coombs, 2021). This is especially true in specific flashpoints like when the first health worker is infected by a patient. Given the importance of changing messaging as the temporal crisis distance increases (Kim and Jin, 2020; Kim, 2022), especially around flashpoints (Dalrymple et al., 2016), we ask:
RQ3a: What are the COVID flashpoints in the CDC public?
Some users will have more of an effect on the discourse on CDC publics over time. To better understand what makes for a more influential messenger over time on the CDC public, we ask:
RQ3b: What are the properties of users most committed to the CDC public?
We adapt the public policy discourse framing analytical framework described in Park and Lee (Park and Lee, 2020)[P.641] to model how Twitter is used as a communication channel to and from the CDC throughout a two-year public health emergency - COVID (see Figure 1). The framework suggests that public policy comprises communication between the state and society (public & private). In this framework, “communicative interaction among political actors operates in a multilayered policy landscape…[to advocate] for ”policy goals (in our case, COVID mitigation policy) and formulating the policy tools (e.g., mask and vaccine mandates) needed to achieve these goals. This communication is mediated through either a one-way or two-way channel to the public. On the other hand, the public will adopt policies and support policy tools depending on their needs and persuasions as stakeholders.
Our study shows that the affordances of technology platforms used in communication need to be taken into account when creating a public for public health crises. We found that indeed, most of the frames in the CDC public were affectively (positive vs. negative) and ideologically (liberal vs. conservative) polarized. The most propagated frames focused on polarizing topics including vaccinations, reopening schools, and arguing about whether the pandemic is a hoax. The mentions network was sparse with strong mutual connections, meaning that echo chambers were formed where some members corresponded with each other while not mentioning the others. Users sharing low-credibility sources were less likely to receive mentions, they were more likely to mention others. This shows that they were more committed to sharing their message in the network. The more low-credibility sources are used by an account, the less likely they are to receive mentions from other accounts. However, we also found that verified accounts, which are more likely to be retweeted, also share low-credibility sources. We suggest new elements to the public policy discourse framing analytical framework based on Twitter affordances and user behavior. We also provide design recommendations for future public policy communication that incorporate emerging technologies, especially generative AI. Finally, we introduce the concept of a crisis messaging journey and show how public health organizations like the CDC need to control messaging about emerging infectious diseases for long time horizons. In Figure 1, we show two main flashpoints, one a sentiment change-point, the other focused on how new variants forced in a change in messaging (more are discussed in §7).
Figure 1. This image shows public policy discourse framing analytical framework adapted from (Park and Lee, 2020). We introduced new elements (in yellow) to the model based on our findings. We show how platform affordances, message and user properties, and platform governance are important factors to consider when communicating public health policy for emerging infectious diseases like COVID. Additionally, we discuss how emerging technologies (in gray), especially generative AI introduce both challenges and design opportunities for crisis organizations like the CDC. Finally, we discuss the effects of temporal crisis distance affects CDC messages over time, especially around flashpoints associated with major changes (e.g., new virus variants). The +/- signs represent how each of these components can make crisis communication more or less effective.
2. Related Work
In this section, we start by reviewing Twitter affordances, namely managing audiences and association with content. Next, we analyze digital publics, focusing on polarizing issues. Finally, we review digital communication in public health crises.
2.1. Twitter affordances and audience management
Technological affordances are defined as the relationship between users and ”their perceptions of environments” (Pavelko and Myrick, [n. d.], P.361). Affordances can explain the influence of different technologies on human behavior (Faraj and Azad, [n. d.]; Leonardi and Vaast, [n. d.]). Twitter has numerous platform-specific affordances that enable people to communicate with each other, associate with content, and provide or receive direct (DeVito et al., [n. d.]) or indirect feedback (Randazzo and Ammari, 2023) due to the searchability and permanence of the content (Leonardi and Vaast, [n. d.]). We will focus on two main affordances: (1) hashtags and (2) interactions including mentions, retweets (Kim and Yoo, 2012; Boyd et al., 2010), likes (McDonald, 2021; Huang et al., 2018; Wang et al., 2016), replies (Kim and Yoo, 2012), and quote-tweets (Garimella et al., 2016; Zade et al., 2024).
2.1.1. Imagined audiences on Twitter
Hashtags are keywords that allow users to search for or find a topic of interest. This allows users to build an imagined audience (Litt, 2012) whom they believe to be interested in a particular topic, “to mark participation, assert individual identity, promote group identity, and support or challenge a frame.” (Stewart et al., 2017) For example, hashtags make issues of interest like Black Lives Matter (#BLM) and #stopasianhate more searchable and accessible, which in turn allows users to reach wider audiences (Tong et al., 2022). Additionally, Huang et al. (Huang et al., 2010) argue that hashtags are conversational tools that motivate people to share their thoughts on the topic of interest.
Another way to build an audience is by using mentions (e.g., @user). When a Tweet mentions a user, only the target user receives a notification, but other users can view and search for the Tweet. In other words, users searching for Tweets made by an account will also see tweets mentioning that account. This means one can expand their audience by targeting an account belonging to a person (e.g., politician or journalist) or organization central to the discourse on a particular topic (McGregor, 2019). Users will typically write negative Tweets mentioning a specific account to send a hostile message rather than start a discussion with the target user (Hemsley et al., 2018) with the hope that a provocative message to an influential node in the network would increase the size of their audience.
Retweets propagate messages throughout the Twitter network, and whether intended or not, serve as endorsements of the original Tweet message (Kim and Yoo, 2012) and are thus associated with the virality of a message (Bonsón et al., 2019). Earlier work trained models using the number of retweets to better understand more influential messages in the network (c.f.,(Jenders et al., 2013; Suh et al., 2010; Bakshy et al., 2011)). Likes also propagate messages to a more targeted audience than retweets (Rathnayake and Suthers, 2017; Sekimoto et al., 2020). Thus, they show the popularity of the message since the user “liking” the Tweet is an implicit approval of the message (Bonsón et al., 2019; Sekimoto et al., 2020).
Responses to Tweets are usually associated with commitment to an issue (Bonsón et al., 2019) as it provides a chance to re-frame the conversation (Zade et al., 2024) by providing a positive or negative view, thus pushing the user’s view of the original message (Garimella et al., 2016). Quote-tweets, much like forwarded messages, not only allow the user to push the original message in a new direction, but also “broadcasts” (Zade et al., 2024) this new contextualized view to networks in which they are embedded (Garimella et al., 2016).
2.1.2. Rich Media on Twitter
While social media was traditionally more text-heavy, social media sites like Twitter, Instagram, and TikTok afford users the capacity to create more rich content - including images and videos. This becomes especially useful on Twitter given the words limits of the platform. Adding rich media allows users to extend their text posts (Jin et al., 2016; Gupta et al., 2013).
Shi et al. show that rich media increases the likelihood that users will propagate messages (e.g., through likes and retweets) (Shi et al., 2018; Jiang et al., 2019) and increase the perceived credibility of the message (Yin and Zhang, 2020; Li et al., 2022). Richer messages also increase the persuasiveness (Zhou et al., 2021b) and the reach (Wang et al., 2023) of COVID-19 misinformation.
2.1.3. Verified Twitter accounts
Verified Twitter accounts are central to the mechanism by which users search for and propagate information on the platform (González-Bailón and De Domenico, 2021; Paul et al., 2019) especially in the event of a crisis like COVID (Zhou et al., 2023). In fact, Twitter shared that it would increase the number of local verified accounts sharing information about COVID (Support, 2020) to provide more trusted information in a timely manner (Lunden, 2020). One example Zhou et al. (Zhou et al., 2023, P.22) provide is Dr. Eric Ding (@DrEricDing) who became a verified account on January 25, 2020. Additionally, self-disclosure by verified accounts about their COVID challenges increased the change of self-disclosure in the network (Du et al., 2022).
However, verified accounts can also become “perceived experts” who spread misinformation about vaccines (Harris et al., 2024) especially in regard to the side-effects of new COVID vaccines like the Oxford/AstraZeneca (Hobbs et al., 2024).
2.2. Digital Publics and Crisis Management
When opinions and knowledge are shared in a communicative space, “an arena in which citizens discuss matters of common concern” (Wessler, 2019), it is referred to as a Public Space (Habermas, 1989). A public can be defined as a space to access information along with an amorphous network of individuals (Van Dijk, 1985). Publics allow for knowledge to be widely shared, making “collective thoughts and action possible.’ (Van Dijk, 1985, P.14). When people have a shared awareness of knowledge, they have a better sense of collective power, while the lack of access to shared knowledge leads to detachment (Van Dijk, 1985, P.14).
In these public spheres, stakeholders come together to talk about issues (c.f., (Dewey, 1927; Rauchfleisch et al., 2021)) defined as “socially constructed matters of public interest around which there is contention” (Stoltenberg, 2024, P.2009). These publics evolve over time around specific topics of interest signified by hashtags that convey the content of the messages as described in §2.1 (Bruns and Burgess, 2011).
This in essence leads to the creation of two types of publics: (1) active publics which are connected to the organization managing a crisis (e.g., local and State agencies following FEMA or State public health institutions following the CDC) (Kinsky et al., 2021); and (2) latent publics: which are formed by people who, while affected by a crisis, might not have control over the response (Dalrymple et al., 2016, P.445) but still engage in dialogue about the crisis in real-time (Bortree and Seltzer, 2009; Seltzer and Mitrook, 2007; Taylor and Perry, 2005; Kim and Grunig, 2011).
2.2.1. Hashtags, mentions, and publics
Hashtags like #BLM also signify affiliation with wider communities that signify their values (Zappavigna and Martin, 2018; Freelon et al., 2016). These hashtags serve as focal points, allowing people to share their feelings through storytelling in what Papacharissi refers to as “affective publics” (Papacharissi, 2016). During crises, people search for hashtags to find “situational updates” that are most important to their crisis (Lachlan et al., 2016). Organizations managing a crisis (e.g., FEMA or the CDC), can create or co-opt hashtags that are followed by a large number of users (Dalrymple et al., 2016).
Using mentions on Twitter is another way of creating digital publics that works by targeting central members of a network (Kountouri and Kollias, 2023). For example, when analyzing the use of mentions on Twitter, addressing specific users with negative or harsh messages can increase the polarization of publics (Kountouri and Kollias, 2023). This affective polarization increases dislike towards the out-group (e.g., Republicans attacking Democrats; (Druckman and Levendusky, 2019; Iyengar et al., 2012, 2019))(Brüggemann and Meyer, 2023).
2.2.2. Framing in Crisis Publics
Affective polarization also leads to polarized framing of messages on social media (Entman, 1993). The reason framing can produce affect in those receiving a message is that the sender (e.g., politicians) “attempt to mobilize [people] by encouraging them to think about…policies along particular lines [like] highlighting certain features of the policy[, specifically,] its relationship to” values that are important to them ((Jacoby, 2000) cited in (Chong and Druckman, 2007, P.106)). In other words, framing allows one to define the problem. Hashtags can be a shortcut to describing evaluative statements on potentially polarizing issues like masking (Lang et al., 2021) If different people have different worldviews, if for example, one group considers public health mitigation mandates as a central public health policy, while others view it as a question of personal freedom, they will be ”talking past each other” (Brüggemann and Meyer, 2023; Giglietto and Lee, 2017), therefore, crisis organizations like the CDC need to craft different messages that satisfy various publics (Sellnow et al., 2009).
Earlier work using sentiment analysis evaluates how different publics react to public policies like mitigation measures (e.g., masking, vaccination and school closures; (Ewing and Vu, 2021; Kwok et al., 2021; Manguri et al., 2020)) or economic support to offset lockdown economic losses (Huang et al., 2022). Indeed, public health interventions have been politically polarizing (Haupt et al., 2021; Valiavska and Smith-Frigerio, 2022), especially on social media sites like Twitter (Valiavska and Smith-Frigerio, 2022), where “Liberate” movement and other groups countered public health measures (e.g. lock-downs) by sharing hashtags like #Liberate. These groups were especially championed by political leaders such as President Donald Trump at the time (Haupt et al., 2021). Earlier work shows that the more a message satisfies different affective publics, the more accepted the response message (Jin, 2010; Utz et al., 2013; Kim and Cameron, 2011).
Organizations like FEMA and the CDC aim to frame their messages in a way that counters politicization and instead, focuses on their missions (Vanessa Murphree and Blevens, 2009). Earlier work suggests that these organizations should consider a crisis from different facets, such as the perspectives of different publics, and find the appropriate framing for each (Stephens et al., 2005).
2.2.3. Misinformation and Distrust
Despite this information shared on social media channels being useful to many, increased communication from public health systems has also given rise to conspiracy theories and distrust. Kou et al. explain that conspiracy theories are a type of rumor. During the Zika virus outbreak, social media served as a source of up-to-date Zika-related news, as well as a hub for conspiracy theories about the virus (Kou et al., 2017). Kou et al. revealed that these theories were the public’s attempt to make sense of the significant shift in society caused by the Zika outbreak (Kou et al., 2017). This coping mechanism quickly becomes convoluted when health authorities do not have complete information about the virus causing the outbreak (Gui et al., 2017). The lack of trust in formal sources of crisis-related information spurs greater conspiracy theories. Those in vulnerable groups tend to experience higher distrust in public health and healthcare systems (Abelson et al., 2009). The HIV/AIDS conspiracy theories, for example, perpetuated the idea that the pandemic was planned by the government to target and harm specific populations (Kou et al., 2017). These theories often have significant consequences. In this example, distrust in the government led to AIDS patients refusing urgent testing and treatment (Casley, 2010). Similarly, images and memes were used to spread misinformation (Wang et al., 2023) and hate (Liao et al., 2023; Vishwamitra et al., 2024; Javed et al., 2023) in COVID mitigation discourse.
2.3. Digital Communication in Public Health Crises
Gui et al. propose including the public in the decision-making process behind crisis communication to reduce the distrust between crisis management organizations like the CDC and the public (Gui et al., 2018). An example of such engagement with the public is the organization of a Hackathon to brainstorm COVID-19 solutions in different social domains held in Germany (Haesler et al., 2021).
Usually however, public health agencies tend to utilize social media as a means of broadcasting crisis management information rather than a two-way communication tool (Graham, 2014). By altering this perspective and examining the public’s sentiment through social media interactions, information gaps and overlooked concerns can be addressed sooner, limiting distrust.
2.3.1. Emerging Infectious Diseases: Rolling uncertainty, risk, and public relations
Fox (Fox, 2012) suggests embracing the “inevitable uncertainty” of medicine where scientists and medical professionals have to contend with “certainty for now” instead of more objective certainty. This uncertainty is significantly higher for emerging infectious diseases (EIDs) like Ebola and COVID because the provision of medical expertise and the touting of existing protocols provides an “illusion of certainty.” (Dalrymple et al., 2016) For example, when the CDC protocols set to prevent the spread of Ebola failed, the CDC tended to emphasize what was already known about the disease (Dalrymple et al., 2016, P.455).
Earlier work shows that when organizations like the CDC emphasize certainty in a chaotic uncertain times, this amplifies the uncertainty that the general public feels, and increases doubts about the institutions as well (Dalrymple et al., 2016).
2.3.2. Crisis and Emergency Risk Communication
The Crisis and Emergency Risk Communication (CERC) model describes how, to better manage uncertainty at the height of crisis, an organization has to engage in two-way communication (Reynolds and Seeger, 2005) where the public can give feedback about the management of a crisis (Renn, 2015). Specifically, public facing organizations like the CDC or FEMA need to engage in what Dozier et al. (Waters and Jamal, 2011) call symetric two-way communication which provides for a mutual understanding between the public and organizations which could lead to changes in their communication strategies (Grunig and Grunig, 2008; Waters and Jamal, 2011). However, earlier work has shown that CDC’s online communication has been mostly top-down (from the CDC to users) instead of bottom-up communication (from users to the CDC; (Gesser-Edelsburg et al., 2014).
In the pre-crisis stage, the CERC model suggests that public health organizations should strive to “alliances and cooperation with agencies, organizations and groups.” (Reynolds and Seeger, 2005, P.52). After the initial crisis event (e.g., a novel virus infects a substantial part of the population), public health organizations are tasked with discussing the frame of disease detection (disease symptoms) and disease prevention (taking preventive measures) (Umphrey, 2003; Higgins et al., 2006). In the maintenance phase of the crisis, public health organizations need to communicate healthcare services (services provided by healthcare system), scientific advances like discovering new evidence about the disease, and lifestyle risk factors (e.g., personal habits associated with the disease) (Ngai et al., 2020). Assuming that there is a resolution to the crisis (e.g, effective vaccine is produced), the organization needs to evaluate the communication experience by engaging with the public (Reynolds and Seeger, 2005).
One major challenge faced by scholars lies in the lack of a mechanism to determine different phases of an outbreak (beginning, middle, and end) since they are not predictable, especially in IEDs (Dalrymple et al., 2016, P.457). We discuss the changes between these fluid phases in the next subsection.
2.3.3. Flashpoints of change in a health crisis
Events like the first transmission of a disease from a patient to a healthcare worker (e.g., Ebola spreading from patient to nurse), serve as “flashpoints of change” (Vanderford et al., 2007) that necessitate real-time “rapid redirection on social media.” (Dalrymple et al., 2016) Any such rapid change is challenged by social media affordances like persistence of earlier data shared by public health organizations (see §2.1).
Earlier work shows that as the temporal crisis distance increases, the sentiment of the crisis publics tends to be more positive, and their response more supporting of the crisis organization (e.g., FEMA or a drug company) (Kim and Jin, 2020; Kim, 2022). Though they did not focus on any crisis publics, Valdez et al. used topic modeling and sentiment analysis to measure public mood changes over a period of four months (starting in January of 2020 and ending in early April of 2020) (Valdez et al., 2020). The use of Twitter peaked around the time of lock-downs and social distancing in March of 2020. While the sentiment of the Tweets was initially positive, it steadily and significantly declined by late March of 2020.
3. Dataset
We used the twarc Python library for data mining via the Twitter API to collect the tweets used in this study. This was possible through the use of an individualized bearer token granted by Twitter. The inclusion criteria for the tweets in this study were tweets that were produced by users in the United States, in English, posted between January 2020 and January 2022, and not retweets. Furthermore, we specifically only included tweets that were sent to the official @CDCgov account, from @CDCgov, used the hashtag ‘#CDC’, tagged @CDCgov, or mentioned the string ‘CDC’. In total, we collected 21,843 tweets sent to @CDCgov, 11,810 from @CDCgov, 20,012 using ‘#CDC’, 53,104 tagging @CDCgov, and 162,295 mentioning the string ‘CDC’. Finally, we collected any other Tweets using the same “Conversation ID” which indicate that they are under the same discussion thread. In total, 275,124 tweets were collected. While we did not collect all retweets, we do have the number of times each tweet is retweeted with a total of 3,331,350 retweets.
4. Methods
In this study, we analyze the discourse centered around the CDC using mixed methods. First, we determined the credibility of sources shared on Twitter in §4.1.4. Next, we explored different themes of the discourse using topic models (§4.1.1) and qualitative analysis (§4.1.2). Finally, to better understand the stakeholders and their interests, we clustered users based on their topics of interest and their behavior on the platform (§4.3.1).
4.1. Features used in data analysis
In this section, we present all the features used to train classifiers, regression, clustering and temporal models we used in our analysis. We start with a description of topics and themes created using topic modeling. We then describe how we measured sentiment, identified low-credibility sources, and finally used platform signals (e.g., retweets) and user properties (e.g., verified user).
4.1.1. Topic Modeling with BERTopic (N=71)
We used the Bidirectional Encoder Representations from Transformers (BERT) for contextualized topic modeling. Due to the limited length of tweets (maximum of 280 characters) and the frequent use of characters like # and @, traditional topic modeling techniques like latent data analysis (LDA) (Blei and Lafferty, 2006; Hays, 1998) tend to produce less intelligible results (Egger and Yu, 2022). Instead, we use a topic modeling method that relies on clustering word embeddings. Word embeddings model the semantic context (Mikolov et al., 2013) of words by mapping corpus terms in semantic space “in which distance represents semantic association” (Griffiths et al., 2007) (cited in (Angelov, 2020)). We used the BERTopic package (Grootendorst, 2022) to cluster BERT embeddings (Reimers and Gurevych, 2019) using Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) (Campello et al., 2013). Similar documents in the corpus will be closer to each other. Each document will also be closer to words semantically closer to it.
Topic Coherence
To find the optimum number of topics for the BERTopic model, we trained 25 different models by changing the minimum cluster size (an HDBSCAN hyperparameter that controls the size of the clusters) by increments of 10 at a time. [^2] https://hdbscan.readthedocs.io/en/latest/parameter_selection.html?highlight=min_cluster_size#selecting-min-cluster-size The minimum cluster size of the first model was set at 15, and the last model was set at 255. The higher the minimum cluster size, the lower the number of clusters and identified topics. To identify the best model, we calculated the coherence score of each model using the Gensim CoherenceModel feature (Řehůřek, nd). Coherence values have been found to be good in approximating human ratings of a topic model “understandability” (Röder et al., 2015). We selected the model with the highest coherence score (0.39), with a minimum cluster size of 45, that produces a topic model with 71 topics.
4.1.2. Qualitative Analysis
In a virtual public square, discourse constructs social contexts (Schiffrin, 1987; Howarth, 2000) as it reflects communication emphasis as expressed by participants (Potter and Wetherell, 1987). When qualitatively analyzing discourse, communication content becomes more than information exchange (i.e., mentions) or diffusion (i.e., retweets) (Matassi and Boczkowski, 2023). Discourse can become a resource to structure knowledge and make sense of a situation (Hardy et al., 2005). Topic models must be interpreted by humans (Chang et al., 2009; Dou et al., 2011) to contextualize them in the social/political environment (Baden et al., 2020; Yarchi et al., 2021). Once the 71 topics with the highest coherence scores were identified, we used qualitative methods to describe each topic and create overarching themes of the issues shared between topic groups as described in Table 15. The following are the steps used in this process:
  1. (1)
    First iteration - describing the top themes: We sampled 100 tweets randomly from each topic to create a general outline of themes that represent frames in our dataset.
  2. (2)
    Selecting top positive, negative, and neutral sentiment comments for each theme: All of the themes had a sentiment distribution with more extreme scores than a normal distribution (Kallner, 2017). Given this distribution, we use log likelihood ratio (LLR) and Tag2Vec to develop top words describing different sentiments associated with each theme. These are presented in Tables 16, 17 and 18. LLR determines whether terms have “topic signatures” that “define the aboutness” (Gupta et al., 2007) of two different sets of words. It was used earlier by Chancelor et al (Chancellor et al., 2018) to determine normative differences between online communities. Similarly, Tag2Vec, using embeddings (similar to the ones discussed in §4.1.1), provides a list of the words closest to a tag (in this case, belonging to a cluster or being tagged as a positive/negative framing of a theme) (Chen et al., 2017). It was used before to determine the linguistic differences between throwaway and pseudonymous users on Reddit (Ammari et al., 2019). We randomly sampled 100 tweets for each category of top theme X sentiment. When there was no sample across any of the sentiment criteria, we sampled another 100 tweets from this theme. This gave us a total of 6,000 tweets.
  3. (3)
    Consensus on themes: After an iterative coding process and weekly meetings to discuss individual code books, we reached consensus on the different themes and the polarization within each (McDonald et al., 2019). We also identified overarching themes as shown in Figure 2.
A note on quoting tweets and identifying Twitter accounts
Fiesler and Proferes (Fiesler and Proferes, 2018) argue that Twitter users do not expect to be quoted verbatim in academic research, even if they are not named. They suggest engaging in Bruckman’s (Bruckman, 2002) recommended levels of user disguise when quoting users in a research study. To protect the privacy of Twitter users, we primarily describe the contents of tweets without direct quotations. We only name public organizations such as the CDC or news organizations and organizations.
Figure 2. Topic groupings into themes
Figure 2 shows the breakdown of topics into five main themes covering a similar discourse around COVID mitigation and politics. §4.3.1.
4.1.3. Sentiment analysis (N=1)
To understand how the sentiments of tweets affected the way that people discussed topics in the CDC public, we used VADER, a sentiment analysis tool that is frequently applied to Twitter text (Elbagir and Yang, 2019) which calculates the valence and polarity for each word (positive, negative, or neutral) along with the percentage of said polarity in a composite score. VADER labels text as having positive sentiment if the composite score is above 0.05, negative sentiment if the score is below -0.05, and neutral sentiment if the score is between -0.05 and 0.05 (Hutto and Gilbert, 2014). We used composite score in regression models, and we used positive, negative, and neutral labels for classifiers (see §4.2).
4.1.4. Credibility of Information Sources (N=1)
We used the Iffy Index of Unreliable Sources, [^3] https://iffy.news/index/ a public-facing dataset of low-credibility sources [^4] https://docs.google.com/spreadsheets/d/1ck1_FZC-97uDLIlvRJDTrGqBk0FuDe9yHkluROgpGS8/gviz/tq?tqx=out:csv&sheet=Iffy-news reviewed by Media Bias/Fact Check, [^5] https://mediabiasfactcheck.com/methodology/ and was used to determine the credibility of medical data about COVID-19 vaccination information on Twitter (Pierri et al., 2023). This dataset contains the following 12 features:
  1. (1)
    Domain: Site address without “https:” and “www.”
    (e.g. google.com instead of https:www.google.com)
  2. (2)
    Site rank: The Alexa rank of a site’s traffic
  3. (3)
    Year online: The year the site was established
  4. (4)
    Name: The site name
  5. (5)
    Factual categorization: Very high, high, mostly factual, mixed, low, very low
  6. (6)
    Bias (least biased/pro-science, right-center/left-center, left/right, questionable/conspiracy pseudoscience)
The Iffy Index only includes sites with a low-credibility rating and categorizes as either Conspiracy/Pseudoscience (CP) or Questionable Source/Fake News (FN).
4.1.5. Twitter Platform Signals (N=9)
As discussed in §2.1, Twitter provides affordances for propagating messages and expanding your audience. We will use the features identified in the section and listed below:
  1. (1)
    Retweet count (RT)
  2. (2)
    Quote Tweet count (QT)
  3. (3)
    Tweet likes (Like)
  4. (4)
    Direct responses to Tweets (Rep)
  5. (5)
    Rich Media: Tweet uses rich data (e.g., pictures or videos)
  6. (6)
    Verified account
Additionally, we add the following features to describe the use of Twitter in the CDC public:
  1. (1)
    Total number of Tweets by user (count): this feature represents the total number of Tweets for a user over the period of data collection within the CDC public. The higher this number, the higher the commitment of the user to this discourse.
  2. (2)
    Tenure in the CDC Twitter publics (date of last Tweet - date of first Tweet). While the Tweet count provides one measure of commitment to the discourse, this is another measure that shows how long the user was engaged in this discussion.
  3. (3)
    FFR: Following/Follower ratio: the higher this ratio, the higher the number of people interested in their content without the target user following many others. If the value is less than one, then the user can be described as a “mass follower” who follows other accounts for the“sole purpose of gaining more users” for themselves (Anger and Kittl, 2011).
4.2. Understanding platform signals and content analysis
To better understand platform signals, we trained four regressions to measure what increases the likelihood of retweet counts, quote tweet counts, replies to tweets, and likes for Tweets (see Tables 2 and 3).
Additionally, we trained seven classifiers to better understand what exemplified verified accounts, Tweets responding to the CDC, Tweets using low-credibility links, Tweets using rich media (e.g., pictures and videos), positive, negative, and neutral Tweets based on Vader sentiment analysis (see §4.1.3). The top classifiers, along with the positive and negatively predictive features are shared in Tables 5 and 6.
4.2.1. Hyperparameter Tuning
When building all of the models identified above, we used 82 features as described below:
  1. (1)
    Twitter topics as defined in §4.1.1 (N=71).
  2. (2)
    Vader sentiment analysis as defined in §4.1.3 (N=1).
  3. (3)
    Low credibility source identifier as defined in §4.1.4 (N=1).
  4. (4)
    Twitter platform signals as defined in §4.1.5 (N=9).
We set an 80/20 random train/test split and used Sklearn’s standard scaler to scale data before training (Pedregosa et al., 2011a). To find the top R2 in our regression models, we used Random Search Cross Validation (with 5-fold cross-validation) (Vishnu et al., 2023) to investigate the performance of Ridge and Lasso regressions. We tuned for $\alpha$ and the solver ([”auto”, ”svd”, ”cholesky”, ”lsqr”, ”sparse_cg”, ”sag”, ”saga”, ”lbfgs”]) in Ridge. We also tuned for $\alpha$ and selection ([”cyclic”, ”random”]) in Lasso regression (Wang et al., 2011).
As for the seven classifiers, we again set an 80/20 random train/test split and used Sklearn’s standard scaler to scale data before training. To find the model with the top Area Under the Curve (AUC) - a common metric that evaluates a model’s fit (Swets, 2014; Fawcett, 2006; Lemeshow and Hosmer Jr, 1982) - we used Grid Search Cross Validation (with 5-fold cross-validation) to find the model with the best AUC. A model is considered acceptable as long as the AUC value is above 0.7. The accuracy, recall, F1, and AUC metrics are reported in average and for the top model in Tables 4, 5, and 6. We trained our models on the following classifiers: (1) logistic regression classifier; (2) Random Forest Classifier; (3) Gradient Boosting Classifier; and (4) Gaussian Classifier. They full list of hyperparameters used in the training process can be found in the Appendix (§Classifier hyperparameters).
4.2.2. Explainabilty
For the Ridge and Lasso Regressions, we used the Regressors package (Haas, [n. d.]) to determine significant coefficients (presented in Tables 2 and 3).
However, given that six of our top classifiers were random forest classifiers, whilst the seventh is a Gradient Boosting Classifier, we applied post-hoc analysis to explain the models. We used SHapely Additive exPlanations (SHAP), a post-hoc explainable AI analysis tool to explain each model (Huang and Huang, 2023; Shoemaker et al., 2023). Shapley values, based on game theory (Lundberg and Lee, 2017), allow us to analyze the effects of each feature on output of the model. In turn, this allowed us to rank top predictors and to show dependence between predictors.
4.2.3. Analyzing mention ties: Exponential Random Graph Models
Exponential Random Graph Models (ERGM) model patterns within a network which can be seen as evidence for ongoing structural processes, which are also influenced by node attributes (Robins and Lusher, 2012). ERGM models “extend the logic of multivariate regression to relational data: the presence or absence of a binary tie in a network is an outcome variable predicted by a vector of independent variables called parameters.” (Keegan et al., 2012)[p.431] A positive parameter increases the likelihood of a tie in the network, while a negative parameter indicates a decreased likelihood of tie formation. The parameter equivalent to the intercept in a regression model is the edges parameter which “reflects the log-likelihood of a network tie appearing entirely by chance” which in turn describes the “network density if no other effects were present.” (Keegan et al., 2012)[p.432]. Another network-specific or “endogenous” ERGM parameter is mutuality, or reciprocity which measures the tendency for a tie from one node to another to be reciprocated in the other direction. A higher mutual value increases the value of “cooperative network ties“ between network nodes (Axelrod and Hamilton, 1981). Exogenous node-level coefficients indicate whether actors with a specific attribute are more or less likely to be connected to alters in the network (Lusher et al., 2013).
The method has been used in HCI and CSCW to describe the dynamics of social networks. Keegan et al. (Keegan et al., 2012) built a bipartite network connecting Wikipedia articles and editors who worked on them, finding that the likelihood of an editor contributing to the Wikipedia article increases with every fatality associated with the incident, and where incidents occurred in OECD countries (Keegan et al., 2012)[p.432]. In the medical field, Srinivasan and Jiang (Srinivasan and Jiang, 2023) created a disease co-occurrence network and showed that chronic diseases increase the likelihood of a tie to COVID diagnosis, thus proving a link between the two.
ERGM parameters
First, we will describe the network we formed to analyze using ERGM. Twitter users can explicitly “mention” other users by including their usernames in the text of the tweet. A user u, can mention another user v, by tweeting “@v, what are you up to today.” We created a directed mentions network based on user mentions composed of 59,304 nodes and 315,495 edges. We then used the Statnet R package (Handcock et al., 2019) to fit the ERGM model on the mentions network using the following parameters: (1) Endogenous parameters: (a) edges; (b) mutuality and (2) exogenous parameters: (a) average theme weights as defined in §4.1.2 given the computational cost of incorporating all 71 topics (Kim, 2025); (b) user sentiments broken by positive, negative, and neutral based on the VADER score Positive, negative, and neutral (see §4.1.3); (c) credibility of sources as presented in 4.1.4; and (d) user-specific features as presented in §4.1.5.
4.3. Exploring discourse in CDC Twitter publics
To identify different publics within the CDC public, we will cluster users into different groups that share topics of discussion and user behavior.
4.3.1. Clustering to find communities
Since our data is not labeled, we used the features presented earlier (topics, sentiment, the properties of the Twitter account) to cluster Twitter users in our dataset. Similar to methods adopted by Hussein et al. (Hussein et al., 2021), we used K-Means clustering on two feature types:
  1. (1)
    Discourse-specific features:
    1. (a)
      Average topic weights discussed by users (see §4.1.1)
    2. (b)
      Average user sentiment when discussing the topics identified in 1(a) (see §4.1.3)
    3. (c)
      Credibility of sources used by Twitter users (see §4.1.4)
  2. (2)
    User-specific features presented in §4.1.5
When using K-Means models, choosing the optimal K number of clusters is of central importance to the quality of the clusters. We used the Yellowbrick Python package (Bengfort et al., 2018) to select the optimal number of clusters using the knee point detection algorithm (Antunes et al., 2018). The optimal k=16 is presented in Figure 3.
Figure 3. Using the elbow method, we identified 16 clusters as being the best number based on distortion error.
4.3.2. Analyzing cluster content
For each cluster, LLR and Tag2Vec were used to determine the signature of the biographies of cluster users. We also used the top topics discussed by each cluster to guide our qualitative analysis. For larger clusters, 100 users were selected (including at least 10% verified users). For clusters with less than 100 users, all users were qualitatively analyzed. The tweets of each user were read along with their biographies after dividing the samples between team members.
As we did for the topics and themes, the team reached consensus on the description of different clusters after regular team meetings following individual coding of each cluster and sharing memos about their findings with the team (McDonald et al., 2019).
4.4. Temporal Analysis
Earlier work also analyzed the temporal dynamics of online discourses and how they are affected by the onset of COVID (Lee et al., 2021) by specifying one breaking point, March 11th, which is when COVID was declared a pandemic. The authors then used the Chow test (Wooldridge, 2019) to measure the significance of the change in topic-level discussions before and after that date. However, given that it is difficult to define the different phases of a health crisis beforehand (see §2.3.3), we describe the change-point detection method we used in §4.4.1.
While change-point detection allows us to describe temporal change at the population level, we also wanted to better understand the temporal patterns of use on a personal level. Similar to the study by Gao et al. (Gao et al., 2021), we analyzed users’ commitment to engagement in the online community over time. We describe our survival analysis in §4.4.2.
4.4.1. Temporal Change-point Detection
We applied the Pruned Exact Linear Time (PELT) change-point detection algorithm (Truong et al., 2020) to identify significant changes in Tweet sentiment and the weight of different topics and their parent themes (as described earlier in §4.1.1).
Change-point detection algorithms find points in a time series where there is a significant change (Killick et al., 2012). We chose the PELT algorithm for change-point detection because it is more conservative in detecting temporal changes, which produces more accurate results (Wambui et al., 2015; Liu et al., 2013). Earlier work adopted this method to analyze Twitter discourse about COVID over time (Valdez et al., 2020).
Once we identified a change-point, we wanted to determine the trend direction (upward or downward trajectory) and significance. To do so, we relied on the Mann-Kendall non-parametric test as presented by Hussain & Mahmud (Hussain and Mahmud, 2019). In figures showing the change-points (e.g., Figures 6, 8, and 10), we used a green spot to identify significant trends, and a blue arrow showing the direction of the trend.
To describe the difference before and after a significant change-point, we used log-likelihood ratio (LLR) to determine “topic signatures” before and after the change-point (Dunning, 1993; Lin and Hovy, 2000) based on the “aboutness” of word lists (Gupta et al., 2007) before and after the change-point (Lee et al., 2021).
4.4.2. Survival Analysis
Following the method used in Gao et al.(Gao et al., 2021), we define user commitment as the duration of engagement in discourse with the CDC in units of participant-month. In our case, the control variables are: (1) retweet count; (2) reply count; (3) like count; (4) quote count; (5) ratio of low-quality links used; (6) posting pictures; and the (7) user being verified. Our independent variables measure engagement in the five main themes identified in Figure 2, as well as average VADER sentiment score for the user.
All continuous variables (other than verified) are standardized using the Sklearn standard scaler (Pedregosa et al., 2011b). We used the lifelines Python package for our analysis (Davidson-Pilon, 2019).
Following the method adopted by Gao et al. (Gao et al., 2021), we used the timestamp for the first Tweet as the starting date for participating in the public sphere. The date of the last Tweet was used as the date of the last activity. Participation is considered censored if the last Tweet is dated from the month the data collection was performed. This is because we do not know when they stop contributing in the public sphere. If a user does not Tweet within three months, we consider them as having dropped out. The model we trained had a concordance of 0.612, which is an acceptable prediction for survival analysis (Zhang et al., 2022b). A summary of the model output is presented in Table 10.
Since the continuous variables are standardized, a hazard ratio (HR) larger than 1 means there is a higher than average likelihood of not surviving (dropping out of the CDC discussions). A hazard ratio less than 1 shows a likelihood of continuing to engage in the CDC public sphere.
5. Study I: Platform signals and message properties
We start this section by analyzing how CDC discourse is affected by Twitter platform affordances (e.g., retweets, likes, etc.) in §5.1.1. Next, we analyze how credibility of sources in a Twitter message, the use of rich media, and the sentiment of that message affect COVID discourse in the CDC public in §5.2. Finally, in §5.3, we analyze how user properties, like being verified or influential in the network (based on FFR), are associated with user behavior in the CDC public.
Summary of findings:
Discussions about vaccines are more likely to receive all kinds of Twitter engagement. Although positive tweets discussed the seriousness of COVID-19, they also shared news about dropping infection levels, especially after the lockdowns. Neutral accounts were more likely to discuss the lack of clarity in the CDC guidelines but avoided talking about mortality rates or engagement in political discourse–both features predictive of negative Tweets. We found that verified accounts, while less likely to Tweet, are more likely to be retweeted (thus the high echo score in Table 1), liked, receive responses, but not quote-tweets, and to use rich media. Tweets using low-credibility sources are also more likely to use rich media and be posted by verified accounts. The more frequently an account uses low-credibility sources, the more likely they are to mention other users.
Table 1. This table compares the CDC, responses to it, verified users, and responses to them, tweets using pictures and responses to them, as well as the use of low-credibility sources and responses to these tweets. Specifically, we focus on the echo and average sentiment for each. Echo is the ratio between retweet/tweets. The closer the value to 1, the less the likelihood of their tweets being retweeted. The higher the echo the higher the number of retweets per Tweet. At times, the properties might overlap. For example, we show how often verified users use images.
Behavior and Affordances Total Unique Echo Sent. Low-cred. (%) Rich (%) Ver. (%)
CDC 2,688,505 11,810 227.65 0.17 0 100 100
Respond to CDC 122,813 21,843 5.623 -0.01 4.16 34.50 6.12
Verified 2,903,733 26,428 109.87 0.10 0.10 75.61 -
Respond to verified 139,058 26,205 5.31 -0.01 3.96 35.07 11.02
Rich media 3,032,625 99,004 30.63 0.01 11.54 - 20.18
Respond to pictures 136,131 27,783 4.90 -0.01 4.28 43.98 8.07
Low credibility 36,008 17,433 2.07 -0.16 - 65.55 5.11
Respond to low credibility 702 468 1.50 -0.07 62.61 61.11 2.78
5.1. Platform affordances, user properties, sentiment and information credibility
We start this section by analyzing how Twitter engagements afforded by the platform like retweets, quote-tweets, likes, mentions, and responses are associated with different themes of COVID discussions. We will specifically focus on direct responses to the CDC and, in turn, direct responses to users from the CDC (as opposed to broadcast messaging).
5.1.1. What is in a retweet and what gets more likes?
As Table 2 shows, discussions for or against the vaccine are more likely to be retweeted and to receive likes. While discussions of lock-downs and re-openings, especially in schools (Topics 38 and 65), were less likely to receive likes, they were retweeted at a higher rate. Similarly, engaging in Topic 1, introducing data showing that COVID is not a hoax, increases the chance of being retweeted while reducing the chance of receiving a like.
The chance of receiving likes and being retweeted is higher when Tweets are posted by verified users and when there is a high number of Tweets in the CDC public. Users with a higher follower-following ratio (FFR) and more direct replies are also more likely to receive likes. However, the more direct replies received by a user, the less likely they are to be retweeted. We discuss direct replies next.
5.1.2. What gets more direct replies and quote-Tweets?
Engaging in discourse for and against the vaccine (Topic 0) increases the chances of receiving a response and being quote-tweeted. The same applies to discussing masking in schools (Topic 26). Both of these topics were polarizing as they pit people who believed that each mitigation tool was important to control COVID spread and others who were against their implementation because (1) they thought that the mitigation strategies are not as effective as advertised; or (2) they thought that forcing these mitigation measures were against their freedom. A related topic early in the pandemic focused on the very presence of COVID as a virus. While some were questioning the magnitude of the effects on society, others were sharing data, especially from the Johns Hopkins dataset, to counter data from the CDC and other official channels that were considered not forthcoming about the effects of the virus. However, Topic 38, focusing on the effects of lock-downs on society, received fewer responses and quote-tweets. Interestingly, this topic received fewer likes as well. However, it did increase the number of retweets.
While being verified is more associated with being retweeted, being a verified user makes it less likely to be quote-tweeted. Furthermore, quote-tweeted users have a lower follower-following ratio (influence in the network) and less Tweets in the CDC public. Next, we focus on analyzing direct responses to the CDC.
Responses to the CDC
As presented in Table 4, the use of rich content was more predictive of responses to the CDC. Much of the rich content included what users referred to as “receipts,” screenshots of earlier CDC messages showing the evolving masking guidelines after vaccinations (Topic 4). Other top topics in responses include the vaccines (Topic 0), the role of the CDC in evictions post-COVID, discussing how COVID is airborne (Topic 36), and comparing the response to COVID with the response to the H1N1 breakout under Obama’s administration (Topic 56).
The framing of these responses went in one of three directions: (1) arguing that the CDC’s guidelines are not strict enough; (2) arguing that the CDC is introducing mandates without the right to do so; (3) the CDC is overreacting since the effects of COVID are overblown (see Figure 20).
Replies to the CDC were less likely to receive direct replies, likes, retweets, quote-tweets, and usually had lower follower-following ratios.
Direct responses from the CDC
We found only 58 direct responses from the official CDC account to 25 accounts based on a question or suggestion. The responses revolved around providing information about the shingles or flu vaccines. A few others were answering questions about the danger of contracting COVID-19 by touching products imported from China. The @CDCgov account answered those accounts by saying there is no danger in handling products from China.
Now that we analyzed responses to and from the CDC, we will next analyze mention co-occurrence networks along with the @CDCgov account.
Table 2. This Table shows top regression results and the top predictors for Retweets ($\alpha=0.31$, solver=svd) and Likes ($\alpha=0.34$, solver=svd). Table shows significant predictors from regression model. **** p¡0.0001; *** p¡0.001; ** p¡0.01; * p¡0.05
Retweets Ridge $R^{2}$=0.90 Likes Ridge $R^{2}$=0.83
Feature Coefficient p-value Feature Coefficient p-value
Topic 0 0.611 **** Topic 0 3.511 ****
Topic 1 0.101 * Topic 1 -0.614 *
Topic 38 0.072 ** Topic 38 -0.432 **
Topic 65 0.045 * Topic 65 -0.23 *
Verified 2.329 * Verified 15.033 ****
like_count 2838.533 **** quote_count -33146.7 ****
quote_count 7534.939 **** reply_count 1189.377 ****
reply_count -1084.07 **** retweet_count 47676.71 ****
count 0.412 **** count 8.882 ****
FFR 17.385 *
Table 3. This Table shows top regression results and the top predictors for responses ($\alpha$=0.31, solver=svd) and quote-tweets ($\alpha$=0.68, solver=cholesky), Quotes-tweets, and Responses. Table shows significant predictors from regression model. **** p¡0.0001; *** p¡0.001; ** p¡0.01; * p¡0.05
Responses Ridge $R^{2}$=0.64 Quote-tweets Ridge $R^{2}$=0.70
Feature Coefficient p-value Feature Coefficient p-value
Topic 0 0.161 **** Topic 0 0.008 ****
Topic 1 -0.06 * Topic 15 -0.008 **
Topic 15 0.036 **** Topic 26 0.011 *
Topic 26 0.004 **** Topic 38 -0.005 *
Topic 38 -0.001 **** Verified -0.032 ****
Verified 0.732 **** reply_count 856.81 ****
like_count 32.959 **** retweet_count 962.706 ****
quote_count 2964.705 **** count -1.813 ****
retweet_count -504.582 **** FFR -0.666 ****
count 10.537 ****
FFR 2.828 ****
Table 4. Predictors for direct replies to the CDC. The table shows the average metrics from RandomSearch and the hyperparameters for the top model. They also show the top positive and negative predictors as determined using SHAP.
Reply to CDC
Average Metrics
Accuracy 0.645
Precision 0.626
Recall 0.767
F1 0.684
AUC 0.716
Top Model (GBC) est=150, max. depth=5, learning rate=0.3
Accuracy 0.752
Precision 0.725
Recall 0.811
F1 0.766
AUC 0.837
Positive Features 0, 4, 27, 36,56, Rich
Negative Features Rep, Like, FFR, RT, 7, 15, 15, 61, QT, 29, 9, 19, 8, 56
Table 5. Part 1 of Classifiers: Performance metrics for Positive, Neutral, Negative, and Reply to CDC classifiers. The table shows the average metrics from RandomSearch and the hyperparameters for the top model. They also show the top positive and negative predictors as determined using SHAP.
Positive Neutral Negative
Average Metrics
Accuracy 0.668 0.681 0.688
Precision 0.667 0.684 0.715
Recall 0.717 0.713 0.593
F1 0.685 0.689 0.633
AUC 0.700 0.719 0.727
Top Model (RF) est=50; mins=5, max=None est=100; mins=5, max=10 est=100; mins=5, max=10
Accuracy 0.984 0.983 0.985
Precision 0.985 0.983 0.984
Recall 0.982 0.983 0.985
F1 0.984 0.983 0.985
AUC 0.999 0.998 0.999
Positive Features RT, 15, QT, FFR, VR, 1 5, FFR, 21, 14, 9, 32, 63 19, 25
Negative Features 25, Rep, 16, 2, 54, 23, 8, 14, 44,29,56 70, Rep, QT, Like FFR, 14, 30, RT, 21
Table 6. Part 2 of Classifiers: Performance metrics for Photo, Low Quality, and Verified classifiers. The table shows the average metrics from RandomSearch and the hyperparameters for the top model. They also show the top positive and negative predictors as determined using SHAP.
Rich Media Low Credibility Verified
Average Metrics
Accuracy 0.719 0.699 0.830
Precision 0.820 0.708 0.838
Recall 0.529 0.754 0.812
F1 0.619 0.718 0.822
AUC 0.756 0.789 0.905
Top Model (RF) est=100; mins=5, max=10 est=100; min=2, max=None est=100; mins=5, max=10
Accuracy 0.996 0.782 0.894
Precision 0.998 0.803 0.876
Recall 0.994 0.748 0.916
F1 0.996 0.775 0.896
AUC 0.995 0.872 0.964
Positive Features RT, FFR, QT, Like, Rep, 5 19, Rich, 7, VR FFR, QT, RT, Like, Rep, 14, Rich
Negative Features VR, 6, 16, 61 Like, Rep, RT, 23, 32 Count, 21
5.1.3. What is in a mention?
Mentioning accounts can be directing a message to an account expecting a response or increasing attention to the message by mentioning an influential account in the network, in this case, the CDC. The co-occurrence mentions network (Nasim and Rajput, 2017) shown in Figure 4 where the top accounts mentioned alongside the @CDCGov account are presented in the following groups:
  1. (1)
    Pharmaceutical companies like @pfizer (specifically focused on the vaccine production)
  2. (2)
    Public health professionals/academics: @DrEricDing (see §2.1.3)
  3. (3)
    Politicians: @VP, @secazar, @joebiden, @govrondesantis, @mikepense, @presssec, @kamalaharris, @statedepartment, @nycgovcuomo,@joebiden, @gop, @statedept, @cdcdirector
  4. (4)
    News corporations and journalists: @foxnews, @jaketapper, @cnn, @abc, @cbsnews, @nbcnews @msnbc, @maddow, @alexberenson (this last account is a verified journalist who happens to share low-credibility information sources; more in §5.3)
  5. (5)
    State and federal agencies: @fema (Federal Emergency Management Agency), @nih, @us_fda (US Food and Drug Administration), @dhsgov (Department of Homeland Security), @hhsgov (health and human services), @cdcmmwr (Morbidity and Mortality Weekly Report), @drnancym_cdc, @HealthyFla (Florida Dept. of Health; see §6.2.1)
The results of our ERGM model, which analyzes the likelihood of a mention tie, are summarized in Table 7. Our edges (intercept) parameter is negative and significant. This means that there is a high ‘cost’ to the creation of a tie in the network (Keegan et al., 2012)[p.432]. In our sparse network, we found a positive “mutual”parameter suggesting a higher likelihood of tie creation associated with reciprocation. In other words, when one user mentions another, they are more likely to mention the original user.
Figure 4. Mention co-occurrence network showing top accounts mentioned along with the CDC account.
When analyzing node characteristics, we found that most of the discussion themes identified in §4.1.1 decreased the likelihood of creating a mention tie. The only theme that increased the likelihood of the creation of a mention tie was the one focused on non-COVID discourse probably due to the high connectivity between those working with opioid addiction patients, pain activists, and veterans (discussed in more detail in §6.3)). We found that out-degree tie formation was more likely for users who posted more using less credible sources. In other words, users who relied on less reputable sources were more likely to mention others while they themselves were less likely to be mentioned. Both negative and positive sentiment decreased the likelihood of the creation of mention ties, while neutral sentiment increased the likelihood. The more the use of rich media in Tweets, the less likely the creation of mention ties. Finally, verified users are less likely to have mention ties with other users.
Table 7. Exponential family random graph model of mentions network. A smaller value means the likelihood of creating a mention tie is smaller and vice-versa. **** p¡0.0001; *** p¡0.001; ** p¡0.01; * p¡0.05
Estimate Std. Error z value
nodeocov.verified -0.062 0.012 -4.979 ***
nodeicov.verified -0.205 0.013 -15.236 **
nodematch.top_theme.Mitigation -0.147 0.007 -20.031 ***
nodematch.top_theme.Non-COVID 1.236 0.032 38.792 ***
nodematch.top_theme.Politics -0.205 0.011 -18.738 ***
nodematch.top_theme.Test -0.267 0.022 -12.250 **
nodeocov.lowq-credibility 0.096 0.024 3.948 ***
nodeicov.lowq-credibility -0.231 0.029 -7.980 ***
nodeicov.photo -0.267 0.008 -32.107 ***
nodematch.sentiment.negative -0.237 0.009 -25.842 ***
nodematch.sentiment.neutral 0.381 0.011 33.583 ***
nodematch.sentiment.positive -0.078 0.008 -9.781 ***
edges -8.896 0.004 -2000.335 ***
mutual 3.483 0.063 55.150 ***
5.2. Sentiment, rich media, and low credibility sources
In this section, we will analyze the qualities of positive, neutral, and negative labeled Tweets, as well as ones labeled with the use of rich media (e.g., images) and low-credibility sources.
5.2.1. Positive Tweets
A higher retweet and quote Tweet count, being verified, and having a higher follower-following ratio are all more positively predictive of positive Tweets. Sharing data sources pointing to the seriousness of COVID (Topic 1), but also to good news about dropping rates of infection (Topic 15).
Positive Tweets are less likely to receive direct responses. They are also less likely to discuss COVID and seasonal flu infection and mortality rates (Topics 16 & 25), the efficacy of masks (Topics 2 and 54), or calling for obedience to the CDC mitigation guidelines in general (Topic 44). They tend to make less mention of political discourse like Trump’s CDC (Topic 8), the political/legislative response to COVID (Topic 29), or comparing the COVID response to H1N1 response under Obama’s administration (Topic 56).
5.2.2. Neutral Tweets
Higher follower-following ratio increases the chance of a neutral Tweet. Neutral tweets tend to receive fewer responses, quote-tweets, and likes.
Sports players and attendance guidelines (Topic 21) and social distancing guidelines (six feet apart; Topic 63) are also positively associated with neutral Tweets. These are often posts by users with high FFR, usually at a local level, like high school and college athletic departments, coaches, and sports commentators who, while not verified, have large followings. Many of the posts simply stated the current requirements for attending sports events (e.g., in-door masking for intramural sports or social distancing requirements).
The lack of clarity in CDC guidelines (Topic 9), the attack and defense of the CDC (Topic 14), and the administrative and legislative response to COVID (Topic 32) are positively predictive of neutral Tweets. While defending/attacking the CDC might sound polarizing, the Tweets themselves can have a more neutral sentiment since they can be nuanced. For example, the author of the Tweet could specify that while they support the CDC’s efforts, they do not think the organization is doing enough to counter the virus. Specifically, they could focus on the fact that the CDC only publishes data when it can have a more positive context – when the COVID infection numbers are decreasing. For reference, Table 1 shows that CDC Tweets are more positive than any other entity, cluster, or theme with an average VADER score of 0.17.
5.2.3. Negative Tweets
Higher retweets and follower-following ratios are less predictive of negative Tweets. The more a Tweet discusses COVID infections and mortality rates (Topic 25) or the opioid pandemic data (Topic 19), the more they are likely to evince a negative sentiment. The first topic can be more negative because it usually focuses on attacking the CDC for reporting mortality numbers that are unexpectedly high or low (note that while COVID infection numbers are heavily discussed in positive Tweets, positive Tweets do not focus on mortality rates). Negative tweets are also less focused on attacking/defending the CDC (Topic 14) and in-person gathering guidelines.
5.2.4. Tweets using rich media
Tweets using rich media are more likely to be retweeted, liked, quote-tweeted, get a reply, and engage with the CDC (Topic 5). Tweets using rich media are less likely to be posted from verified accounts, or to discuss lack of up-to-date COVID testing data (Topic 6), seasonal flu mortality rates (Topic 16) or music and stress (Topic 61).
5.2.5. Tweets using low credibility sources
The use of rich media and verified accounts are more predictive of the use of low credibility sources. The focus on discussing the opioid pandemic (Topic 19) and COVID mortality rates (Topic 7) are also positively predictive of low credibility sources.
Tweets using low credibility sources are less likely to be liked, retweeted, get responses, or engage in discussions regarding legislative and administrative responses to COVID (Topic 32) or Salmonella outbreaks (Topic 23).
5.3. Verified accounts and accounts with high follower-following ratios
Verified accounts are more likely to be retweeted, quote-tweeted, liked, get direct responses, use rich media and engage in attacking or defending the CDC (Topic 14). Verified accounts are less likely to Tweet in the CDC public (Tweet count) and less likely to discuss Sports players and attendance guidelines (Topic 21). While verified accounts engaged with less Tweets than other users, their messages were retweeted at a higher rate which is in-line with the high echo value for verified accounts indicated in Table 1. It is also interesting to note that responses to verified accounts in general, and the CDC in particular were around 5, which means that the user received 5 retweets for an engagement with a verified account or the CDC.
While the use of low-credibility sources is not predictive of verified accounts, verified accounts are predictive of the use of low credibility sources in a Tweet. In other words, some verified accounts used low credibility sources in their engagement. One example is @alexberenson, a journalist who pushed back against most of the CDC’s mitigation strategies. Berenson, with a verified journalist account, used several low-credibility sources and was mentioned along with the @CDCGov account with high frequency (see §5.1.3) which would expand his audience further.
Accounts with high FFR were more likely to receive likes and responses. They were more likely to engage in positive or neutral messaging and use rich media in their Tweets.
6. Study II: Discourse Themes, Interest Publics, and Stakeholder Groups
In this section, we answer RQ2 by analyzing framing used by different publics and how they are polarized, and we identify the main stakeholders engaging in each public.
Using the top 10 hashtags and the co-occurrence network of hashtags that are associated with #CDC (Türker and Sulak, 2018) (Figure 5), we found the following main areas of focus:
  1. (1)
    Mitigation:
    1. (a)
      Masking: #maskup, #masks, #wearamask
    2. (b)
      Vaccines: #getvaccinated, #vaccines,, #covidvaccine
    3. (c)
      Lock-downs: #quarantine, #socialdistancing, #stayhome
  2. (2)
    Science and medical expertise: #NIH, #WHO, #FDA, #DrFauci, Healthcare Common Procedure Coding System (#HCPCS): focusing on medical insurance
  3. (3)
    Political (Trump) and well-known public health figures (Dr. Fauci). #Florida is also the top-tenth hashtag due to the political fights between the Florida state government’s handling of the disease and the CDC guidelines (discussed in more detail in §6.2.1): #MAGA, #whitehouse, #trumpliesamericansdie, #trumpisaliar, #trumpvirus
Table 15 shows these top themes, along with their echo (total retweets/total number of tweets by user), the ratio of tweets by verified users, tweets with rich media, and low-credibility tweets. As shown in the table, all of the themes have a negative kurtosis, or wide distribution with more extreme scores than those in a normal distribution. This means that the discussions were polarized (Kallner, 2017). In this section, we analyze the mostly polarized discussions under each theme, and we introduce the different stakeholders and their respective positive, neutral, and negative communication frames.
Summary of findings:
We found that there are three main areas of discourse in the CDC public: (1) Mitigation guidelines: introducing the different CDC mitigation guidelines and how changes to them over time (not recommending masking for all at the beginning, then recommending it, then saying the vaccine will negate the need for masking, until the Omicron variant took hold); (2) COVID politics: how the CDC was not ready for the pandemic, and how the roll-out of the vaccine for the population was considered good progress after losing trust in the CDC early in the pandemic; and (3) Non-COVID discourse: while issues like Salmonella outbreaks and vaping dangers become important as COVID takes hold, some other conditions like opioid addiction continue to be public health issues now compounded by lock-downs and other COVID complications.
Table 8. Cluster top topics Part I
# Cluster Name Topics and Weights (%) RT avg
1 COVID racial disparities COVID racial disparities (73), Voices for and against vaccines (1), COVID testing kits accuracy & availability (1), Using COVID data to counter idea of hoax (1), Hopkins COVID data vs. CDC data (1) 7.78
2 Mitigation policy discussion Using COVID data to counter idea of hoax (5), Mentioning and responding to CDC (5), Mask wearing guidelines (4), Jokes about CDC COVID management (3) 10.85
3 Partisan discussion of COVID response Legislative & administrative response to COVID (25), Trump’s CDC (8), Threat of politicizing CDC science (5), Mentioning & responding to CDC (5), Attacking and defending the CDC (5) 1.41
4 Guideline changes with Omicron Variant Omicron variant changing masking guidelines (78), Voices for & against vaccines (3), Hopkins COVID data vs. CDC data (1), Six feet apart guidelines (1), Air travel guidelines (1) 10.49
5 CDC Voices for & against vaccination (11), Using COVID data to counter idea of hoax (4), Flu vaccinations (3), Salmonella outbreaks (2), Risks to school teachers and students (2) 26K
6 Pro-mitigation Parody of CDC masking guidelines (8), Voices for and against vaccinations (2), Mask wearing guidelines (2), Calls for following CDC mitigation guidelines to control COVID spread (1) 1.53
7 No clarity in CDC guidelines Lack of clarity in CDC guidelines (49), Calls for following CDC mitigation guidelines to control COVID spread (17), Mentioning and responding to CDC (1), Voices for & against vaccination (1), Attacking & defending CDC (1) 0.29
8 Obesity and COVID Guidelines for obese patients (79), Voices for & against vaccination (2), Delta variant changing masking guidelines (1), Mentioning & responding to CDC (1), Hopkins COVID data vs. CDC data (1) 0.84
9 Gun violence Gun violence (80), Voices for & against vaccination (2), Jokes about CDC’s COVID management (1), COVID racial disparities (1), Delta variant changing masking guidelines (1) 6.3
10 Food-borne diseases outbreak Salmonella outbreak (71), Mentioning & responding to CDC (1), Lack of clarity in CDC guidelines (1), Voices for and against vaccination (1), Mask wearing guidelines (1) 2.43
11 Vaccinating pregnant women Vaccinating pregnant women (72), Using COVID data to counter idea of hoax (2), Mentioning & responding to CDC (2), Voices for & against vaccination (1), Attacking and defending the CDC (1) 9.03
12 School sports & community events Voices for & against vaccination (42), Sports players & attendance guidelines (4), In-person gathering guidelines (3), Social distancing re-opening guidelines (3), Florida COVID infection & mortality numbers (2) 4.91
Table 9. Cluster top topics Part II
# Cluster Name Topics and Weights (%) RT avg
13 Air travel guideline changes Air travel guidelines (77), Using COVID data to counter idea of hoax (1), Social distancing re-opening guidelines (1), CDC COVID infection and mortality rates (1) 1.24
14 Mask wearing guidelines Facial hair and masking guidelines (80), Voices for & against vaccination (2), COVID testing kits accuracy & availability issues (1), Evolving masking guidelines with vaccinations (1), Calls for following CDC mitigation guidelines to control CDC spread (1) 0.76
15 School reopening dynamics Risks to school teachers & students (75), Voices for & against vaccination (1), Vaccinating teachers (1), Neglecting COVID-related adverse childhood experiences (1), Mentioning and responding to CDC (1) 4.4
16 Quarantine dynamics Quarantine guidelines (77), Using COVID data to counter the idea of a hoax (1) Voices for & against vaccination (1), Lack of up-to-date COVID testing data (1), Dr. Fauci’s role in managing COVID (1) 1.22
6.1. CDC mitigation guidelines
In this theme, discourse focused on different mitigation policies the CDC has called for. These include the following: (1) masks; (2) vaccines; (3) social distancing, and (4) COVID test-trace. Discussions about these topics were polarized between those who thought CDC guidelines were not stringent enough and those who did not believe the CDC had the right to mandate any guidelines. They revolved around discussing and clarifying CDC mitigation guidelines with the introduction of new variables, such as variants of interest (e.g., the Delta and Omicron variants), as well as the evolving understanding of the efficacy of vaccines.
Figure 5. The figure on the left shows the co-occurrence of hashtags with the #CDC. The figure on the right shows the top hashtags in the discourse. We collapsed similar hashtags together. For example, ‘#coronavirus’, ‘#covid19’, ‘#covid-19’, ‘#sars’, ‘#coronaviruspandemic’, ‘#pandemic’, etc. were collapsed to simply to ‘#COVID-19’. ‘#CDC’ and ‘#cdc’ were collapsed to ‘#CDC’. The hashtags ‘#vaccine’, ‘#GetVaccinated’, ‘#vaccinated’, and ‘#VaccinesWork’ were collapsed to ‘#Vaccine’. Likewise, ‘#Trump’, ‘#TrumpVirus’, ‘#TrumpLiesAmericansDie’ were collapsed to ‘#Trump’. Lastly, ‘#Masks’, ‘#WearAMask’, ‘#MaskUp’, ‘#masks’, and ‘#mask’ were collapsed to ‘#Masks’.
6.1.1. Mitigation guidelines in the social context
Discourse about the validity of CDC guidelines and their effectiveness was a major area of discussion. Cluster 2 (see Table 9), one of the bigger clusters (with about 40,000 Twitter users), which contained almost 2,000 verified users, focused on discussing how COVID is not a hoax. Many of those leading this charge were verified public health professionals (N=23) who explained why CDC mitigation guidelines (e.g., masking and social distancing) are important to control the spread of the disease. Others (N=17) were identified as reporters, anchors, and medical journalists who reported on the initially evolving mitigation guidelines presented by the CDC. However, some users who self-identified as conservative and/or Trump supporters (N=5) questioned the veracity of the pandemic, and many considered it a hoax and used concepts like “plandemic” to explain it away. Those views were sometimes in response to verified public health accounts. Others mocked anyone who ever masked outdoors, even at the height of COVID surges. The frustrations expressed by anti-masking users were couched in the idea of freedom - a word that often appeared in Twitter user descriptions in Cluster 2. Another frame pushed by most users in cluster 8, most of them self-identifying as politically conservative (N=4), question the negative health outcomes from COVID itself. They instead argue that underlying conditions, specifically obesity, are to blame for high death rates. In other words, obesity killed people who happened to have COVID.
Cluster 14 focuses on guidelines for where masks need to be worn (e.g., spaces like grocery stores, hospitals, etc.) and how to wear masks correctly, especially referencing facial hair for men and makeup for women (cosmetology is one of the identifying words in Table 12). The efficacy of the surgical masks is brought into question. This discourse is about N95/KN95 masks and how they are more efficacious than other masks, especially with newer variants. Several public health professionals criticized the CDC guidelines about cloth and N95 masks. Specifically, they were arguing for mandating N95 masks in hospitals and for vulnerable populations. They were also frustrated with the scarcity of N95 masks and other protective gear in hospital settings. To make their point, the opponents of masking mandates posted earlier CDC guidelines that did not recommend masks along with the current recommendations that say otherwise. This is one of the earlier examples of the lack of clarity in CDC guidelines. Three of the users relied on screenshots of earlier CDC guidelines to argue against new, usually stricter, mitigation efforts (e.g., you should wear a mask even if you are not symptomatic).
In Cluster 12, these concerns were localized to school district sports activities and how they can be reopened safely for families, teachers, and student-athletes. While some parents were pushing for a return to ”normal,” medical professionals pointed out that COVID infections of youth athletes can have long-term effects on them. They cautioned against reopening sporting events without adequate mitigation, including COVID testing. Many accounts included high school and college athletic departments (N=11), coaches and heads of athletic teams (N=5), and school administrators (N=2). A focus on relaying the latest mitigation guidelines by the CDC for sporting events was evident in neutral Tweets as discussed in §5.2.2.
Positive Communication Frame: Wearing a mask is important to protect yourself and others
Neutral Communication Frame: KN-95 masks give a higher level of protection than cloth masks
Negative Communication Frame: Forcing people to wear a mask goes against our freedoms
6.1.2. As the pandemic evolves, guideline clarity devolves
Cluster 7 is focused on the idea that CDC guidelines are not clear. The discourse in this cluster was mostly pro-mitigation and leaned on the liberal side politically. For example, in Table 12, we can see that words like #ANTIFA, #BlackLivesMatter, etc. were indicative of the biographies in this cluster. The focus on professional status was also important in this cluster (see words like scholar, chief medical resident etc.). Medical professionals (N=17) wanted the CDC to issue simpler guidelines and not change them in response to subtle changes in the number of COVID infections. Many also pointed out that the CDC should use other metrics besides death rates because the medical and scientific community still has little understanding of long-term effects of COVID infections. The change in quarantine guidelines was critiqued by vascular and pediatric medical doctors in Cluster 16. Specifically, they were concerned about the shortening of the quarantine period from 14 days to five days. This issue was also picked up by different stakeholders in Cluster 13. While reporters were presenting the CDC guideline changes, epidemiologists (N=4) and other healthcare professionals (N=8) as well as medical institutions (N=5) argued against this change. They believed five days did not provide enough protection against spreading COVID to others. However, this was countered by other stakeholders, including the Air Line Pilots Association and travelers, who professed to have lost patience with all the new travel guidelines. In a related way, many posters were excited about CDC guidelines that do not require tests or masking on airplanes or indoors if fully vaccinated. Some medical professionals (N=5) and advocates (N=4) still encouraged masking to protect themselves and others. There was disappointment in the new guidelines and worry among essential workers about un-vaccinated individuals potentially exposing them to the virus. As masking guidelines were changing, questions arose about the efficacy of the vaccine in stopping infections. These questions emerged especially as the vaccine became less effective after the Delta and Omicron variants (Cluster 4). Some of the discourse touched on reintroducing masking, especially at peak COVID infection points (e.g., the Omicron wave in late 2021). Another point of friction focused on the changes that came with re-opening schools. Both are discussed in more detail in §7.1.
Positive Communication Frame: While the vaccine will not protect you from the Omicron variant, it still protects from severe illness
Neutral Communication Frame: Because the new variants are different from the wild-type COVID, vaccines are less effective against them
Negative Communication Frame: Why did the CDC say that the vaccine was over 90% effective if it does not protect against the new variants?
6.1.3. School management post-COVID
Cluster 15 is focused on school openings. The discussions in this cluster are split between those pushing for opening schools back up (N=10), three of whom self-identified as parent advocates. One of the parent advocates argue school closures have negatively affected the managing of adverse childhood experiences (ACE) since many of the mandatory reporters (e.g., social workers, teachers, school councilors, etc.) do not have contact with children. Teachers (N=7), superintendents (N=3), and teacher union leaders (N=1) called for teachers to have a higher priority on the vaccination list to potentially allow for faster school re-openings, which made many parents happy. However, since the vaccines were not approved for children at the time, those same accounts were worried about COVID being transmitted between students. Therefore, these accounts called for masking in schools regardless of vaccination status. At the same time, many of the advocate parents that were introduced earlier also perceived masking as unsafe and unfair to children. Indeed, we found that school re-openings and mitigation guidelines can be polarizing and engaging more retweets (§5.1.1) and responses (§5.1.2).
Positive Communication Frame: Masking, which protects both teachers and students, is more important as new variants appear
Neutral Communication Frame: These are the precautions that players and parents need to take for high school sports events to continue
Negative Communication Frame: Closing schools and forcing students to mask are both affecting students negatively
6.2. Politics
School management was not the only area of the pandemic response that was politicized. In this section, we will show how Twitter users focused on the lack of governmental preparation that was evident in the lack of COVID tests. We follow this with a wider discussion of government response (or lack thereof) to the pandemic. Finally, we focus on racial disparities in the healthcare system in general and through the COVID prism in particular.
6.2.1. Test-trace and mortality rates
The lack of COVID test kits, especially in the early days of the pandemic, is central to this theme. As seen in Table 12, the accounts of public health professionals (N=5), medical research institutions (N=2), and self-identified liberal politicians (N=3) engaged in critiquing, and at times directly attacking, the Trump administration’s management of testing. The main issue they focused on was that accurate reflection of infection counts is important for safeguarding others, especially in healthcare settings (e.g., #nurselife). They used hashtags like #wherearethetests, #liesaredeadly, and #incompetentinchief. The focus on the lack of testing and questioning mortality rates was associated with negative Tweets as discussed in §5.2.3. There was a specific focus on the COVID testing strategy in Florida where the @HealthyFla (Florida Dept. of Health) determined that only symptomatic people would be tested for COVID as opposed to testing a larger sample of the population to assess whether lock-downs are needed. Governor Desantis’ account (@govrondesantis) is frequently mentioned in this discussion as many point to his influence in this COVID testing strategy. An example of rich media messages about the topic can be seen in Figure 22. Given how politicized the testing theme has been, we will explore politicization of the discourse further next.
Positive Communication Frame: While there is currently a lack of COVID testing kits, supply chains are being created
Neutral Communication Frame: It is currently difficult to measure the number of COVID deaths because of the lack of tests
Negative Communication Frame: Not having an adequate supply of tests is a direct consequence of bad planning by the CDC and the administration
6.2.2. Politicization of COVID response
In Cluster 3, the discourse focused on the threat of the politicization of COVID. At the outset of the pandemic, the CDC was frequently labeled as “Trump’s CDC” because the pandemic budget was slashed just in time for the COVID pandemic, which deeply affected the preparedness of the American public health system. Many of the users in this cluster used signals like #VoteBluenoMatterWho, #Resist, #TheResistance, #BlueWave, and #BidenHarris. Three users self-identified as housing advocates of a moratorium on evictions, which was eventually invoked by the CDC.
There is a more positive view of the vaccination roll-out plan as more people receive the vaccine and the focus moves from lock-downs and pressure on the healthcare system. However, public health researchers (N=4) and reporters (N=6) addressed how COVID had a “racial bias,” as its effects on racial minorities were significantly worse without any response to this health outcome bias from state institutions. We focus on racial disparity in more detail next.
Positive Communication Frame: Thank you to the CDC, to the scientists, and to government officials who created the policies that allowed me to get the vaccine so I can go back to my normal life!
Neutral Communication Frame: CDC advising insurance companies that PrEP and similar medications should be fully covered under current insurance plans! This is imperative for HIV prevention.
Negative Communication Frame: People are dying and COVID is spreading even faster! Why is Food and Drug Administration (FDA) pushing its meeting to certify the vaccine?
6.2.3. Racial disparities in a pandemic
Cluster 1 focused on health equity for Black, Latino, and Chicano populations by self-proclaimed advocates from these communities (N=7). These include non-profit organizations (N=3), such as the American Descendants of Slavery (ADOS) and the National Association for the Advancement of Colored People (NAACP), academic institutions (N=11), and faculty (N=9) who researched different areas of racial disparities, such as lack of access to vaccines or their higher exposure to the disease as front-line workers. Similarly, medical institutions and scholars advocated for more equitable access to vaccines for pregnant women of racial minorities using hashtags like #HealthEquity and #VaccinesSaveLives. This framing focused on how minorities, especially because they serve in front-line jobs that were deemed essential, bore the brunt of COVID infections, and they suffered higher death rates and other negative health outcomes. The discussions called for universal health coverage and better access to resources (e.g., PCR testing), especially for members of minority groups. While discourse focusing on the intersection of race and healthcare decreased overall in the CDC public, cluster 1 continued its focus on Topic 40 (COVID racial disparities) throughout the period of the study (see §7.1).
The NAACP is thankful for the equitable roll-out throughout the state. Learn where you can get your vaccine [through this link] to protect yourself and your family.
Neutral Communication Frame: The CDC says racism is a threat to public health outcomes. This is a good step to addressing health inequities.
Negative Communication Frame: While local health departments are tracking COVID infections and deaths without grouping by race we cannot alleviate structural racism in public health.
6.3. Non-COVID CDC guidelines
While most of the discourse in our dataset has focused on COVID, there were other issues that the CDC continued to manage at the time. For example, veterans and pain advocates discussed the need for relaxing CDC rules for opioids, especially for veterans who might have particular needs based on their injuries. They also noted that isolation from the COVID shutdown has added to the pain and mental health challenges faced by veterans. Some of the accounts identified as pain advocates (N=7), many of whom were veterans (N=11) connecting with others who share similar challenges, highlight how their pain was especially exacerbated by the COVID lockdown and pressures that affected their peer support group activities.
As Table 12 shows, words like MAGA, #Liberal, #GunReformNow, #GunlawsSaveLives, and gun-related trauma are all representative of the way users in Cluster 9, which focuses on the polarizing issue of gun control, described themselves. While gun reform was the focus, many self-identified libertarians (N=3) framed their freedom and belief in the Second Amendment as their rejection of CDC mandates. On the other hand, advocate organizations like Moms Demand Action, which are interested in public safety through gun law reform, framed CDC guidelines as central to community safety. One of the less polarized clusters was Cluster 10, which revolved around food-borne diseases (e.g., Salmonella) about which the CDC continued to report.
Positive Communication Frame: Remembering my sister who passed one year ago today that happens to be #opioidawarenessday. You can find Narcan and social support at our center if you need help [address]
Neutral Communication Frame: The CDC is providing data showing that fentanyl is the #1 cause of death in Americans age 18-45. Plans need to be made to address this issue.
Negative Communication Frame: The #OpioidHystera that has been led by the CDC and has caused severe pain patients to be treated like criminals.
7. Study III: Tweeting through the pandemic
In Study II (§6), we found that periods of policy change introduced more uncertainty to the public. For example, clusters 2, 4, 7 (focusing on the lack of clarity in CDC guidelines), 13, 14, 15, and 16 all focused on policy change, like removing the need to mask after receiving the second COVID vaccine shot, a policy that was then changed with the Omicron variant wave (cluster 4). Given that it is important to understand flashpoints in crisis management (see §2.3.3), we used temporal models including change-points and survival analysis (detailed in §4.4)) to analyze: (1) main flashpoints over more than two-years of the pandemic (§7.1 and §7.2); and (2) behavior that increased the chance they would remain engaged in the CDC public (§7.3).
Summary of findings:
In §7.1, we describe three COVID crisis phases based on change-point that tracked early discussions about the seriousness of COVID, the introduction of mitigation guilines, and changing them with new variants. We then showed how the pattern of rich media use coincided with changes in CDC messaging about mitigation (e.g., the return to masking after vaccinations due to new variants). Finally, we showed how the use of rich media, being verified, sharing low-credibility sources, and being retweeted all contributed to long-term engagement on the CDC public.
7.1. Evolving discourse on CDC-related publics: three phases of the COVID crisis
In this section, we report our findings on pandemic phases using change-points as detailed in section 4.4.1. In our findings, we show that there there three main phases: (1) hoaxes and early pandemic effects; (2) lock-downs and masks; and (3) post-vaccine variants and evolving guidelines.
7.1.1. Beginning Phase: Hoaxes and early pandemic effects
Early on in the pandemic, discussions about the race disparities in healthcare (see Figure 7) and communication about the dangers of vaping (Figure 6 both dropped significantly and remained low throughout the period of the study. This shows that in the early days of 2020, COVID is becoming more of an emergency with LLR (see 4.4 for details) values showing that Topics 62 (social distancing guidelines), 59 (parody of CDC guidelines), 57 (facial hair and mask guidelines), and 34 (news reflecting the role of the CDC in managing the pandemic) become more prevalent after the first change-point.
While the racial disparities theme showed a change-point trending downwards early in the pandemic, we have identified a cluster focused on race challenges in the pandemic, described in more detail in §6.2.3. To determine if there were any change-points specific to the racial disparity cluster, we repeated the PELT analysis, specifically for the Cluster 1 subset. This is presented in Figure 7. As you can see, there is a continuous focus on Topic 40: COVID racial disparities throughout most the duration of this analysis (note that the duration for the cluster is not the same as the rest of the dataset because it covers a subset of 121 users). This shows that flashpoints detected on the community level could be different than those that represent a specific set of stakeholders forming an interest public of their own - in this case, the healthcare equity public.
The next change-point is for Topic 1: Using COVID data to counter the idea that it is a hoax. We found that there are two change points, one on the 21st of January 2020 and the other is on 31st 2020. Both these change-points show a downward and significant trend suggesting that this topic peaked between two dates. Topic 48 (Hydroxychloroquine and Ivermectin controversy) was one of the main topics associated with the first change-point. As the name suggests, this topic focused on the veracity (or lack thereof), of the use of these two drugs to treat COVID. While some Tweets were sharing information showing their lack of efficacy, less traditional links were also shared about the need to use these medications that are available on the market.
Both change points for Topic 1 increased the discussion of Topic 62 (social distancing guidelines). Topics 59 (parody of CDC guidelines), 57 (facial hair and mask guidelines), and 34 (news reflecting the role of the CDC in managing the pandemic) are associated with both change-points. These topics focus on the guidelines shared by the CDC as the effects of the COVID emergency become clearer. While some Tweets focused their framing on taking COVID more seriously given the seriousness of COVID,
Finally, one of the change-points in the Testing Theme (see Figure 8) focused on the lack of available COVID tests and the low accuracy of the available test (see details in §6.2.1).
Figure 6. This figure shows change-points for Topic 1 (using data to counter idea of hoax), and Topic 45 (vaping dangers), the mask them. As evident in the figure, all three have significant change-points early in the pandemic. The mask theme has a second change-point discussed in the second phase. Green dots show significant trends and blue arrows show trend directions.
7.1.2. Middle Phase: Lock-downs and masks
The introduction of lock-downs and the call for masking guidelines represent middle part of the COVID crisis. The discourse here started with a discussion of the lack of direction from the CDC (parodies of CDC guidelines).
After the second change-point by late March for Topic 1, the discourse shifted more towards Topic 30: in-person gathering guidelines. This shows how after the first lock-down order was issued in the US on March 15th of 2020, the focus became more on managing the risks associated with COVID, rather than discussions around whether COVID is a hoax, or making light of changing CDC guidelines. However, some posts took a different approach framing the CDC’s guidance as being more draconian. One of the areas of discussion was the suggestion of masking outdoors or double masking (e.g., wearing two masks). Others criticized what they saw as the constantly changing guidelines coming out of the CDC. This is especially true for masking.
Discussions about masking peaked between March 31st 2020 and June 14th 2021 (see Figure 6). Discourse around the first change-point early in the pandemic is focused on guidelines on wearing masks and the CDC’s role in managing the pandemic (Topics 30, 57, 62, 34, and 59). After the first change-point, many Tweets focused on the fact that the CDC changed its position on masking from the early days of the crisis when the guidance was that only healthcare workers and anyone showing symptoms should mask while later on, masking was suggested for all in-doors activities. One example comparing earlier guidelines is presented in Figure 19. We discuss the second masking change-point in §7.1.3.
Figure 7. While there is an early change-point showing that Race discourse decreased early on in the pandemic, COVID Racial Disparities (Topic 40) is heavily discussed in Cluster 1, showing that change-points vary across different publics.
7.1.3. End Phase: Post-vaccine variants and evolving guidelines
Towards the tail of the period covered in this study, we found that change-points are detected in connection to: (1) vaccination; and (2) new variants. The first change-point is about Topic 0: Voices for and against vaccinations. The significant change-point was detected on December 26th 2021. This period of time coincided with the hight of the Omicron peak (as shown in Figure 9). Using the LLR signatures for discourse post change-point, we found that the two top topics signifying post-change-point discourse are Topics 47 (Omicron variant changing masking guidelines) and 31 (Delta variant changing masking guidelines). Both these topics focused on discussing if the new variants reduce the efficacy of the vaccine, and by reducing the efficacy of the vaccine, require the re-introduction of masking guidelines. Mitigation guidelines the focus of Topics 30 (in-person gatherings guidelines), 38 (effects of closures and re-openings on the economy), and 62 (social distancing guidelines), all of which were in flux because new variants reduced the efficacy of the COVID vaccines.
Similarly, discourse after the Mask Theme’s second change-point on June 14th (see Figure 6) focused on the evolving guidelines due to new variants (Topic 47: Omicron variant changing masking guidelines, Topic 31: Delta variant changing masking guidelines, Topic 62: social distancing guidelines). The CDC’s messaging about the changes associated with the new variants as they reduce the effectiveness of the vaccines are seen by many as confusing. They point to images like Figures 17 and18 where, while the CDC is reflecting how the new variants are less effective with the vaccine, and thus might require wearing masks especially in-doors, the claim is still that the vaccine is still effective against hospitalizations (see Figure 15).
A similar pattern is true for the School Management Post-COVID theme. The early change-point focused on discussions around social distancing and masking as the first school-year was being planned after the school closures early in the pandemic. The second change-point focused on the evolving guidelines and challenges associated with new variants. An example of this is shown in Figure 13 where the CDC image shows how the Delta variant is changing CDC guidelines for schools specifically because the vaccine is not as effective as it is to the wild-type virus.
Figure 8. This figure shows change-points in the medium phase and later phases, showing Topic 0 (voices for and against vaccination), COVID Test Theme, and the School Management Theme. Green dots show significant trends and blue arrows show trend directions.
Figure 9. COVID cases in Delta and Omicron waves
7.2. Change-points in CDC Tweet frequency, frequency of rich media content, and overall sentiment
When looking at the use of rich media in Tweets, we found two change-points, one early in pandemic on the 25th of February 2020, and then again on December 4th 2020. Both these change-points are significant with an upward trend. the LLR showing the top topics after the first change-point shows that pictures were first being used to relay detailed information about mitigation guidelines (Topic 62: social distancing guidelines, Topic 57: Facial hair and masking guidelines, Topic 48: Hydroxychloroquine and Ivermectin controversy).
The second change-point with rich media increases focus on Topics 4, 31, and 47 (Delta and Omicron variants changing masking guidelines) as well as evolving masking guidelines with vaccination where the CDC and other public health accounts were calling on the public to mask again since the new variants render the vaccine less effective. These change-points are in-line with the findings presented in §7.1. The change-points associated with “receipts,” usually screenshots of earlier Tweets from the CDC that might have provided different guidelines.
As the PELT timeseries in the middle of Figure 10 shows, Tweets by the CDC dropped to a minimum between February 25th 2020 and late May 2020. The trend after the change-points is positive, indicating that Tweets from the CDC were on the rise. However, between the two change-points, the topical focus was on criticizing the CDC. Specifically, making a parody of CDC masking guidelines (Topic 59) and the lack of their engagement on issues like the Hydroxychloroquine and Ivermectin controversy (Topic 48). This shows that when there was a vacuum in messaging from the CDC about the medical emergency, other members of the CDC public filled that message void by critiquing the lack of engagement.
The mean sentiment in the CDC public was positive until a change-point on November 1st 2020. The significant trend of after the change-point is negative. Topic 29 (Political/legislative COVID response) is the top LLR value after this change-point showing that there is a higher focus on the politics of COVID - from the mandatory introduction of vaccines, to reducing the number of lock-downs, to discussing whether government should be introducing a support package. One of the reasons for this might be the continued cycle of lock-downs and re-openings by this point. Specifically, there was significant push-back against in-person gathering guidelines for thanksgiving (see Figure 21). While this came later, Topic 47 (Omicron variant changing masking guidelines) also had a high LLR after this change-point, showing that reverting back to more stringent mitigation guidelines lowered the average sentiment in the CDC public.
Figure 10. This figure showing change-points for rich media, Tweets by the CDC, and overall Vader sentiment. Green dots show significant trends and blue arrows show trend directions
7.3. User commitment over time
In this section, we use the hazard ratio of the variables reported in Table 10 to determine the likelihood of their continued engagement in the CDC public. Any hazard ratio less than 1 increases the chance of their continued presence. The method is detailed in §4.4.2.
Being a verified user increases the likelihood of continued engagement (by 18%). Using rich media at a rate that is 1 standard deviation from average picture use is the most predictive of continuing engagement at 36% higher likelihood. The retweet variable shows that the those users who get retweets that are one standard deviation higher than average were 7% more likely to continue engaging in this discourse. Those users who use low-credibility sources of information at a rate one standard deviation higher than average, is 6% more likely to continue their engagement.
Interestingly, the higher the like count, the lower the likelihood of continual engagement (at 8%). Engaging in any of the main themes at a level 1 standard deviation higher than average reduce the likelihood of continued engagement (except for School and Vaccine themes which did not have any significant effects). When engaging in more positive sentiment at a value 1 standard deviation higher than average, the lower the likelihood of continued engagement in the community (6%).
These findings suggest that being a verified user, and using pictures in their Tweets, and being retweeted increases the chances of their continued engagement. While at a lower rates, the higher the use of low-quality information sources, and negative sentiment in their Tweets also increases the likelihood of continued engagement. On the other hand, focusing on any of the themes reduces the likelihood of continued engagement.
Table 10. Survival analysis predicting how long members continue to Tweet in the CDC public sphere. **** p¡0.0001; *** p¡0.001; ** p¡0.01; * p¡0.05.
Covariate Type Hazard Ratio p
Retweet Count Platform signal 0.93 *
Low-credibility Behavior 0.94 ****
Rich Media Platform signal 0.64 ****
Verified Platform signal 0.82 ****
Like Count Platform signal 1.08 *
Masks Theme 1.03 ****
Mitigation Theme 1.02 ****
Non-COVID Theme 1.03 ****
Politics Theme 1.02 ****
Test Theme 1.02 ****
VADER Sentiment 1.06 ****
8. Discussion
At the top of this paper, we introduced the disaster of the smallpox outbreak in Milwaukee 130 years ago. Then, as now, public health crisis communication is quite political. The main difference that current-day public health institutions have to grapple with is a more complex sociotechnical communication ecology.
We argue that platform affordances and emerging technologies, especially in generative AI, are all important to incorporate into the public policy discourse framing analytical framework (Figure 1). We first discuss the lack of crisis message clarity in a continuously evolving pandemic enmeshed in polarized political framing seen through the lens of Twitter affordances in §8.1. Next, we propose design recommendations based on our findings, specifically, a CDC AI assistant in §8.2. Finally, we present policy implications for our findings to support the CDC and other public health organizations in managing online communication for emerging infectious diseases in §8.3.
8.1. Crisis messaging journeys, platform affordances, and verified users
In this section, we reflect on our research questions to better understand CDC communication through the lens of Twitter affordances and message properties (§8.1.1), polarized publics (§8.1.2), and message journeys through pandemic flashpoints (§8.1.3).
8.1.1. RQ1: Engagements, sentiments, and Twitter affordances
The importance and low levels of two-way communication
There were more than 21 thousand unique Tweets responding to the CDC directly, averaging five retweets for each Tweet, using rich media like “receipts” which included earlier messages from the CDC (see §5.1.2). However, the CDC itself only responded directly 58 times to 25 unique accounts. This result is in line with earlier findings about how the CDC and related organizations used Twitter as top-down communication channel in H1N1 (Gesser-Edelsburg et al., 2014) and Ebola (Dalrymple et al., 2016). Two-way communication is central to the CERC model (Reynolds and Seeger, 2005) because it increases mutual respect and the potential to incorporate the views of the public into organizational communication practices (Dozier et al., 2013). We propose design recommendations to support more bottom-up communication in the future in §8.2.1. Public understanding of science is made in collaborative communication between scientists and members of the public (Felt and Davies, 2020). However, this communication is especially susceptible to being taken out of context and/or used in misinformation (Allcott et al., 2019) which we discuss next.
Low-credibility sources, verified accounts, and rich media
We found that users who use low-credibility sources are more likely to continue posting on the CDC public and to mention other users in the network, while using the rich media which can increase the persuasiveness of misinformation (Zhou et al., 2021a). While low-credibility Tweets themselves might not be propagated at a high rate, they might be when created and shared by verified accounts. In fact, as we described in §5.1.3, we found that one of the verified journalists (@alexberenson) has used less credible sources in a number of his Tweets which countered the CDC’s mitigation messages (e.g., lock-downs, masking, and vaccinations).
Our findings show that the use of rich media increases the propagation of Tweets (see Table 1) and that Tweets with low-credibility sources were more likely to use rich media (see 5.2.5). Therefore, the development of more complex generative text2image and text2video models (Zhao et al., 2023) which allow users to quickly produce rich media, presents another future moderation challenge. While Tweets using less credible sources were not heavily propagated in our dataset (see §5.1.1), it is difficult to determine if this is caused by any platform-level algorithmic moderation intervention (c.f., (Duskin, 2024; Sharevski et al., 2022)). Similar content, though not using the same sources, could be propagated at higher rates. This calls for more flexible and adaptive moderation models that could find and counter this content. We describe one in §8.2.2.
Sentiment valence not always representative of message framing
While earlier work took negative sentiment to be a proxy for opposing public health policy (e.g., (Unlu et al., 2023)), our work shows that negative sentiment might indicate that the user does not think the public policy went far enough (e.g., reducing the number of quarantine days; see §5.2.3). Similarly, we described in §5.2.2 how posts with neutral sentiments can provide nuanced critiques of current policy while also forwarding public health messages in local publics (e.g., high-school event guidelines). We continue our analysis of affective publics (Valiavska and Smith-Frigerio, 2022) next.
8.1.2. RQ2: Polarized affective publics and public health
In §6, we presented a number of publics along with examples of affectively polarized frames (Tables 9, 12, 16, and 17) along with an array of stakeholders (e.g., teachers, public health professionals, parents, racial equity advocates) who in many cases espouse an us vs. them engagement (i.e., #MAGA vs. #VoteBlueNoMatterWho) while framing central issues like vaccination in two different planes - public safety vs. personal freedom (i.e., #vaccinessavelives vs. #vaccinerisk) (Brüggemann and Meyer, 2023).
This polarization was also evident in our network analysis. We found that the high likelihood of reciprocity coupled with the sparsity of the mentions network centered around the CDC (§5.1.3) suggests the creation of filter bubbles (Barta and Andalibi, 2021). When analyzing propagation, we found that discussing polarizing issues like vaccinations and lock-downs increased the chance of retweets and likes which might also provide another avenue for the creation of filter bubbles (see §5.1.1).
All these findings suggest that users on the CDC public see filtered content based on who they follow, the hashtags they search for, and suggestions made by recommendation systems (filtering algorithms). In turn, the behaviors place users in bubbles of curated information (i.e., filter bubbles;(Barta and Andalibi, 2021; Sundar and Lee, 2022). While earlier work shows that filter bubbles can polarize people by distorting their realities in their echo chambers (Pariser, 2011), marginalized communities can benefit from being in such filter bubbles where they can find other users with similar experiences (Randazzo and Ammari, 2023). For example, cluster 1, which focuses on racial health disparities of COVID on marginalized communities, might be a beneficial space for organizations (e.g., ADOS), researchers, and others who might be interested in this issue. If a new entrant to the network mentions the CDC, or uses #CDC, and asks about racial disparities in health outcomes in the context of COVID, then the recommender system can introduce them to central nodes in cluster 1. We describe a relevant design recommendation in §8.2.4.
Given its mission as a public health institution, the CDC needs to reach out to people “where they are” (Mathews et al., 2020), we will introduce a design recommendation for reaching out to different issue publics in §8.2.4. We argue that by finding and engaging with publics of interest, the CDC can focus on crafting narrative messages that can fit the needs of each community (Sandell et al., 2013; Alamir et al., 2024).
8.1.3. RQ3: Managing friction in crisis messaging journeys from one flashpoint to the next
In §2.3.3, we discussed how it was difficult to gauge different phases of a public health crisis, especially when it is an emerging infectious disease, like COVID. Dalrymple (Dalrymple et al., 2016) argues that there is a need for a rapid change of communication strategy around these flashpoints, and that, due to social media affordances (§2.1), especially persistence and forwardability, changing communication strategies can be difficult.
In Study III (§7), we used change-point detection and an analysis of Tweets before and after those events to describe two phases that hinged on a change in guidelines based on new data and a better understanding of the virus as well as the introduction of new factors like vaccines and new virus variants. In both of these cases, responses to the CDC included “receipts” of earlier CDC messaging. This might be explained by the perceived lack of honesty about the CDC’s uncertainty vis-a-vis an emerging infectious disease (Holmes, 2008). In fact, Holmes et al. (J. Holmes et al., 2009) suggested that focusing on “certainty for now” could earn public trust more than conveying certainty and then being proven wrong about IEDs in future flashpoints.
This is in line with earlier work that shows how data friction (challenges of sharing and interpreting data) (Edwards et al., 2011) increases with the distance between message creators and re-users (Borgman and Groth, 2025) and with the duration of time from the original message to the time of reuse (Bates et al., 2016; Borgman and Groth, 2025). The different “ways in which [data] are …modified to fit different uses as they travel across space, time and social situations” are defined as data journeys (Leonelli, 2020). We introduce the concept of crisis messaging journeys which describes how crisis messages initially shared by the CDC change based on context, purpose (e.g., opposing new mitigation guidelines), and time elapsing from the initial CDC message (Gregory et al., 2024).
Guidelines shared by the CDC at any time during the pandemic might change later given new information gleaned about the EID (e.g., while the vaccine is still effective against hospitalization, people can still be infected with new variants). While scientists understand that “scientific knowledge is always provisional, always open to revision,” (Edwards, 2013, P.427) public engagement with scientific knowledge (i.e., members of the public using “receipts” of earlier guidelines to argue against the veracity of updated guidelines) can “represent ideological and political strategies [that] represent a new form of friction…inclusion friction [which can] promote confusion, suspicion,” and misinformation (Edwards, 2013, ibid).
There are three ways to reduce inclusion friction: (1) using more cautious language (e.g., employing caveats in public science communication (Stocking and Holstein, 1993, P.200)) when presenting guidelines signifying “this is what we know now, rather than this is what we know”(Gregory et al., 2024); (2) linking the original message either by quote-tweeting it to explicitly show the evolution of science and/or annotating (Zheng and Ma, 2022) the original message using an affordance like community notes (Kankham and and, 2024); and (3) pre-exposure (Lewandowsky et al., 2012, P.116) or pre-bunking (Barman and Conlan, 2024) where the audience is presented with an alternative narrative (Lewandowsky et al., 2012, P.117), thus filling any gaps which might result from changing messages in a flashpoint. This would include messaging to explain the reasons for new guidelines by comparing them with their original messages and explaining how the difference was brought about by new medical information. A mechanism to detect flashpoints in real-time is presented in §8.2.3.
Earlier work has shown that if uncertainty claims are made early by scientists in their publics science communication, they are seen as being frank as they “openly admit incomplete knowledge.” (Zehr, 2012, P.9) This is imperative to the success of the crisis organization because the publics’ audiences are more “likely to reject organizational messages during a crisis situation that are inconsistent with past actions” (Stephens et al., 2005, P.399) which would negatively effect crisis communication around flashpoints.
8.2. Design Recommendation: The CDC AI Assistant
Sanders et al.(Sanders et al., 2022) introduce a generative response model based on earlier public health messaging. Indeed, while social media interactions are usually between human actors, new generative AI models like ChatGPT are introducing non-human agents (Quéré et al., 2025; Aubin Le Quéré et al., 2024; Schroeder et al., 2024) which could play a bigger role in exchange of health-related information. Endacott and Leonardi (Endacott and Leonardi, 2022) build on this idea through the conceptualization of artificially intelligent communication technologies (AICTs), defined as “goal-directed, computational agents that use AI to make decisions about communication and communicate on someone’s behalf.” New Generative AI models introduced on social media sites like Facebook are interacting with users of the platform. Meta AI on Facebook, for example, can respond to questions through messenger (Meta, 2025). This is an example of advanced AICTs, or Generative assistants (Park et al., 2023) which have an “architecture that handles retrieval where past experience is dynamically updated at each step and mixed [the] current context.” Extending the design of current offerings (i.e., Meta AI), in this section, we present the CDC AI assistant, which is designed to support the CDC staff based on three triggers: (1) direct response to the @CDCGov account; (2) suspected mis/disinformation (a flag that could be raised by other users or using automated moderation tools); and (3) temporal change-point detection.
Though they can perform helpful tasks, introducing advanced AICTs reduces the organization’s “degree of control over the communication environment” (Park et al., 2023). That is why we need a human-in-the-loop to work on verification of triggers (e.g., is this actually misinformation) and, where necessary, editing of messages to be published by the CDC AI assistant (Wu et al., 2022).
8.2.1. Tailored responses to responses tagging the CDC account
The CDC AI assistant can be trained on earlier messages shared by the CDC account on Twitter as well as content produced on other online accounts. After fine-tuning a generative AI on earlier CDC text, it can reproduce information shared earlier by the CDC on Twitter or available on the CDC website. This can be done in response to the first trigger when the CDC is responded to directly. However, given the sensitivity of messages from the CDC, we will model the assistant after the CoAuthor AI assistant developed in (Lee et al., 2022) where the assistant provided a sounding board to authors. More specifically, given the text of the response to the CDC account, the assistant can present the CDC staff (i.e., the human-in-the-loop) with suggestions based on earlier text and their context (Zhang et al., 2023). To help the CDC staff further, the AI assistant can explain the suggestions using chain of thought (CoT) represented by a step-by-step breakdown of the reasoning (Wei et al., 2022). After the human-in-the-loop drafts the final message, they can save it to an archive of responses along with the original response to the CDC and any annotations they might find helpful. These responses can be used to retrain the AI for future engagements.
8.2.2. Detecting and responding to dis/misinformation
While social media platforms can introduce system-level moderation tools to detect misinformation, these applications are usually not trusted by social media users who would rather use tools that are produced by organizations that do not have a profit motive (Jhaver et al., 2023). Additionally, as crises evolve, the nature of misinformation evolves quickly as exemplified by the COVID “infodemic.” (c.f.,(Wang et al., 2023; Hobbs et al., 2024)) That is why relying on the wisdom of “crowds of ordinary users” to flag dis/misinformation by accounting for “content facticity, user intent, and perceived harm” has proven to be a more effective moderation policy (Bozarth et al., 2023). In our case, users might trigger the CDC AI assistant by mentioning @CDCGov and using a hashtag like mis/disinformation that would then again be annotated and responded to by CDC staff as introduced above. As a substantial repertoire of mis/disinformation content is built up, the assistant can be trained to respond to messages that repeat misinformation even if worded differently by searching for similar text. In addition to community moderation flags, the assistant can also be trained on annotated mis/disinformation datasets (c.f., (Hayawi et al., 2022)) or by flagging low-credibility sources as we did in §2.1. This would allow the assistant to find and counter more disinformation (Sahu et al., 2021).
8.2.3. Real-time flashpoint detection
We analyzed change-points based on content in §7.1 and based on message content and user properties in §7.2. Given the importance of identifying flashpoints and changing communication strategies accordingly (Dalrymple et al., 2016), automating the detection of change-points can help the CDC deal with flashpoints as they occur. Based on a system presented in (Lyu et al., 2024), we first apply a baseline change-point data set from earlier interactions (as presented in §7), which are then incorporated into a Bayesian model. This first phase provided one of two outputs: either the change is a candidate for a change-point or it is not. If it is, a PELT model is applied at set time durations (e.g., every week or month). When detecting a change-point, manual inspection is still important to verify that indeed there is a change after the flashpoint. The manual coder can confirm that indeed, this is a flashpoint, remove it, or if not clear, set it as pending to request annotation by another human-in-the loop. This will allow the CDC to change its communication strategy.
8.2.4. Engaging with different publics and helping users find communities
Given how we described the importance of providing different narratives for different polarized publics (see §8.1.2), creating a multi-agent system would allow the CDC to engage in a more sustained communication process with more users by allowing different generative agents to focus on different publics (Hadfi and Ito, 2022) represented by the clusters we identified in §6. For example, we would build a CDC AI assistant focused on racial bias in healthcare. That agent would engage most with the community in cluster 1 by providing updates in relation to racial bias (e.g., new data released) or responding to messages in the cluster if they are not answered in a specified amount of time (e.g., 12 hours).
The assistant can also suggest accounts at the center of cluster 1 for users interested in the primary topics discussed in the cluster especially if they have not engaged with them before. Given that we are using generative AI, the assistant can also explain why they are being suggested user(s) based on their interests (Lubos, 2024; Di Palma, 2023). For example, the assistant could message a user saying “given your interest in healthcare equity, you might want to follow [list of users in cluster 1].”
8.3. Policy Recommendation: Preparing for the long-now of health crises
Most studies of crisis informatics focus on the response of crisis organization (e.g., the CDC) during early phases of a crisis (Soden and Palen, 2018). However, even in early 2025, on the eve of the fifth anniversary of COVID, around 40% of the population still think the pandemic is “not over,” that they are not back to living their “normal” lives, and are worried about the long-lasting effects of COVID (e.g., long COVID) and of the next health crises (e.g., bird flu) (Gallup, 2025). This shows that there is a need to understand the long-term communication response needs for public health crises like COVID.
Continued engagement with issue publics, especially those with well-defined stakeholders and constituencies (e.g., racial equity and pain advocates), and regional publics (e.g., local help groups providing mutual aid (Soden and Owen, 2021; Randazzo et al., 2025)) can prove especially important in the future should there be any engagement with local populations and their constituencies. Earlier work on local mutual aid groups showed that while they started by addressing immediate needs (e.g., access to food) resulting from the early COVID surge, these same groups advocated for rent control a few months into the crisis (similar to the rent control advocates identified in §6.2.2). Additionally, as mitigation guidelines started being relaxed, the needs of immunocompromised, persons with special needs (Zhang et al., 2022a; Joy et al., 2024; Mack et al., 2022a), chronically ill (Mack et al., 2022b), and long COVID (Karra et al., 2023) populations become paramount.
#CDCChats like the ones presented in (Dalrymple et al., 2016) can be a start, however, the CDC might want to use different platforms depending on population needs. For example, recent work found that many uncertain parents relied on WhatsApp connections to make better sense of COVID (Joy et al., 2024). The design recommendation presented in §8.2 can support the CDC with long-term engagements with different stakeholders. Additionally, the use of data annotated by the human-in-the-loop described above can create more contextualized data. In doing so, Twitter user engagements can provide citizen sensors (Goodchild, 2007) where citizens can “create and share information, forming vast sensing that allow for information in certain topics to be collected, stored, mined, and analyzed in a rapid manner.” (Huang et al., 2022, P.2).
Since earlier work shows how online and offline publics overlap in their crisis concerns (Randazzo et al., 2025), the CDC can model signals from specific regions of the country over time to determine policy changes in real time. This can be done by combining methods introduced by Wen et al. (Wen et al., 2022) who discussed how news topics changed over time in response to policy changes and the work done in Whata & Chimedza (Whata and Chimedza, 2021) who used a dataset of government measures (e.g., lock-downs) and the Google COVID-19 Community Mobility Reports dataset which measures population density in several public space categories (e.g., pharmacies, parks, etc.) over time to determine the effects of government policies in different locales.
Reflecting on the public policy discourse framing analytical framework adapted from (Park and Lee, 2020) which we present in the introduction (Figure 1), we introduce three new components to the model, which we argue should be considered by crisis-management organizations as they engage with the public:
  1. (1)
    Social media affordances: what messages are propagated most and how networks of interactions are formed.
  2. (2)
    Message and user properties: sentiment valence of the message, whether it uses low-credibility sources and/or rich media, and whether it is shared by verified Twitter accounts are all important to take into account in crisis communication.
  3. (3)
    The crisis message journey throughout the “long-now” of public health crises as new flashpoints resulting from new scientific findings (e.g., new variants) and changes on the ground (e.g., vaccine roll-out) are important to manage as they arise.
Additionally, we argue that it is important to understand the different issue, affective, and regional publics (e.g., pro-mitigation or racial health equity clusters). Finally, given the impact of generative AI on how people create content and engage with others, crisis organizations should consider employing these technologies to support their communication strategies while also developing measures to counter their effects on crisis communication.
9. limitations and future work
While our findings show that image use increases the chance of message propagation (see §5.1.1), our work does not analyze the image contents of the tweets over time. Given that images are potent ways to communicate messages, it is important to understand how they factor into public health crisis communication. Future work can use critical discourse analysis methods to cluster and analyze the change in framing in visual content over time (Walters et al., ). Additionally, while our study analyzes a substantial amount of data, we have not engaged directly with users to better understand how they viewed the discourse around COVID in the CDC arena. Therefore, future work should build on Pine et al’s (Pine et al., 2021) analysis by conducting further interviews with users while incorporating not only the way users consumed information in healthcare publics, but how they contributed to the conversation, or are currently transported to new online communities.
10. Conclusion
This study explored the role of Twitter as a prominent “virtual public square” for public health communication during the COVID-19 pandemic. We analyzed 275,124 tweets from, with, to, and about the CDC. We found that as knowledge about COVID changes in different flashpoints, the crisis message journey is affected by publics’ responses comparing the latest changes with earlier guidelines by showing “receipts.” This is associated with polarization around topics like masking and vaccinations. We propose design changes to produce more inclusive virtual public squares and productive public health policy discussions during public health crises, particularly for marginalized communities.
Appendices
Classifier hyperparameters
Hyperparameter tuning was completed using sklearn package’s Random Search Cross Validation class (Pedregosa et al., 2011a). Table 11 shows the details of the hyperparameters used with each classifier.
Table 11. This table shows the different runs of classifications. RF: Random Forest classifier; GBC: Gradient Boost Classifier; LR: Logistic Regression classifier; GNB: Gaussian Naive Bayes. n_estimates: Number of estimates is a hyperparameter used to determine the number trees used in model. max_depth: Maximum depth is the hyperparameter that specifies the level of complexity in a random forest or gradient boosting classifier. The learning rate hyperparameter determines the contribution of each tree to the final prediction in gradient boosting classifiers. The C and penalty hyperparameters determine regularization for logistic regression classifiers. The solver is the algorithm used for optimization
Model n_estimators max_depth min_samples learning_rate C penalty solver
RF 50.0 NaN 2.0 NaN NaN NaN NaN
RF 50.0 NaN 5.0 NaN NaN NaN NaN
RF 50.0 10.0 2.0 NaN NaN NaN NaN
RF 50.0 10.0 5.0 NaN NaN NaN NaN
RF 100.0 NaN 2.0 NaN NaN NaN NaN
RF 100.0 NaN 5.0 NaN NaN NaN NaN
RF 100.0 10.0 2.0 NaN NaN NaN NaN
RF 100.0 10.0 5.0 NaN NaN NaN NaN
GBC 50.0 3.0 NaN 0.1 NaN NaN NaN
GBC 50.0 5.0 NaN 0.1 NaN NaN NaN
GBC 50.0 3.0 NaN 0.2 NaN NaN NaN
GBC 50.0 5.0 NaN 0.2 NaN NaN NaN
GBC 100.0 3.0 NaN 0.1 NaN NaN NaN
GBC 100.0 5.0 NaN 0.1 NaN NaN NaN
GBC 100.0 3.0 NaN 0.2 NaN NaN NaN
GBC 100.0 5.0 NaN 0.2 NaN NaN NaN
LR NaN NaN NaN NaN 0.1 l2 lbfgs
LR NaN NaN NaN NaN 1.0 l2 lbfgs
Cluster users
Tables 12, 13, and 14 show the number of users, including verified users, along with the top biography terms that signify the cluster users for each of the clusters described in §6.
Table 12. Cluster user counts, verified users, and top biography terms
# Cluster Name Users Top desc doc2vec Top desc LLR
1 COVID racial disparities 121 (13) @georgetown, displaced, self-proclaimed, care, oncology, #healthequity, alliance, wellness, Latinx, #Harmreduction, chicano, #Goodtrouble, ADOS law, latino, black, program, medicine, POTUS, community, gender, socially, racist, violence, getvaccinated, epidemiologist, care, feminist, adjunct, diversity, researcher
2 Mitigation policy discussion 39749 (1,932) first amendment,#ANTIFA, #BuildTheWall, #ClimateAction, #CONSERVATIVE, #RightMatters, #AllLivesMatter, #Metoo, #MAGA, Supporter, #Truthmatters freedoms, dweller, proponent, bioethics, thedemocrats, Johns Hopkins University, Come get them
3 Partisan discussion of COVID response 4918(93) #patriot, #bidenharris2020,#independent, GoBucks, #StrongerTogether, #Science, Realist, #TeamPelosi, Latter-Day, #BlueWave, #VoteBlueNoMatterWho retired, trump, resist, donald, steelernation, detest, theresistance, equalrights, hamiltonmusical, election, angry
4 Guideline changes with Omicron Variant 41(6) Educator, @Harvard, Educated, #books, Forbes, @ClemsonUni, lecturer, mgmt, Epidemiology, Strategist, CEO, @HowardU njspotlightnews, action10news, cichealth, byline, columbiamsph, kennedy_school, columbiamed, newscaster, nyuniversity,
5 CDC 1(1) NA privacy, centers, cdc, updates, source, control, prevention, daily, disease, safety, official, health
6 Pro-mitigation 91(1) #Geek, OPINIONS, Makeup, Thinker, #DubNation, #RedSox, #Braves long distance relationship, nonconformist, hospitalityindustry, card, truecrimejunkie,
7 No clarity in CDC guidelines 1280(40) scholar, #ANTIFA, hacker, #SaveOurDemocracy, #BlackTransLivesMatter fanboys, relax, wtvc, dogfather, chiefmedicalresident, finanacialprofessional
Table 13. Cluster user counts, verified users, and top biography terms
# Cluster Name Users Top desc doc2vec Top desc LLR
8 Obesity and COVID 87(2) Kamala, Streamer, hcaker, #research, conservative, stupid, obesity, ASMBS, nutrition, dietitian, lbs, biostatician, diabetes, investigating, sugar,
9 Gun violence 76(4) MAGA, vaccinated, Montanan, #Riseup, #GunReformNow, #Liberal securitywarriors, gunlawssavelives, endoflifeoptions, stopliars, votingrightsforall, trauma, momsdemand, 2nd
10 Food-borne diseases outbreak 278(11) Meat, Pastry, Trash bidenapocalypse, material, consumers,
11 Vaccinating pregnant women 70(12) cweight, news10, childbirth, alabamastateu, newberrylibrary, iumedschool, mayocliniccv, umnmedschool, mildly @sloan_kettering, @EmoryUniversity, @News_8, #HealthEquity, #VaccinesSaveLives, Neuroscientist
12 School sports & community events 7760(584) accredited, supporter reporter, anchor, story, fox4kc, hiv, endorsements, uclaathletics, school, community, pediatric, comorbidity
13 Air travel guideline changes 209(23) Air Line Pilots Association, vacation, cbsaustin, thedailybeast @News_8, @NY1, Nuggets, Toronto, Correspondent, @sloan_kettering, @UCF, @HarvardMed, Epidemiology
14 Mask wearing guidelines 80(6) FAU, Nuggets, @HofstraU cosmetology, #itsthesmallthings, bloggers, linesman, medtechs, imresidency, nyuwagner
15 School reopening dynamics 997(53) daughter, fall, accredited, childhood, counselor, curriculum, practitioner, athletic, vascular teacher, schools, educator, learner, education, superintendent, principal, mom, district, business owner, association
16 Quarantine dynamics 134(8) feet, tired, meaning rights, vol_baseball, vamosgalaxy, annmediasports, childhoodcancerawareness,oxy_football
13 Air travel guideline changes 209(23) Air Line Pilots Association, vacation, cbsaustin, thedailybeast @News_8, @NY1, Nuggets, Toronto, Correspondent, @sloan_kettering, @UCF, @HarvardMed, Epidemiology
Table 14. Cluster user counts, verified users, and top biography terms continued
# Cluster Name Users Top desc doc2vec Top desc LLR
14 Mask wearing guidelines 80(6) FAU, Nuggets, @HofstraU cosmetology, #itsthesmallthings, bloggers, linesman, medtechs, imresidency, nyuwagner
15 School reopening dynamics 997(53) daughter, fall, accredited, childhood, counselor, curriculum, practitioner, athletic, vascular teacher, schools, educator, learner, education, superintendent, principal, mom, district, business owner, association
16 Quarantine dynamics 134(8) feet, tired, meaning rights, vol_baseball, vamosgalaxy, annmediasports, childhoodcancerawareness,oxy_football
Polarization in topics
In §6, we discussed the polarization between and within different clusters. For example, cluster 6 represents those who are pro-mitigation. However, cluster 2 includes stakeholders pushing more stricter mitigation guideline and those who think current CDC guidelines are too strict. Table 15 shows the way different topics were the echoed (retweets per tweet) for different themes. It also shows the % of low-credibility, rich media, and verified in each theme. Figures 11 and 12 show the sentiment distribution for each of the themes. Tables 16, 17 shows the top signal words for the positive and negative sentiments for each of the themes. Table 18 shows signature words for neutral text. Not all themes have neutral labeled text and therefore, not all themes are represented in the table.
Table 15. For each of the top themes, we list the total number of tweets, echo, sentiment distribution, low credibility count, picture count, and the number of tweets by verified users. *Sub-themes under the main CDC mitigation guidelines theme. ** Sub-themes under the Politics theme.
Theme Unique Echo Sent. Low-cred. (%) Rich (%) Ver. (%) Kurtosis
Mitigation 108,340 13.93 0.07 4.54 36.67 $1$1.66
Vaccination* 28,567 11.91 0.05 6.13 41.26 16.24 -1.69
Masks* 20,678 36.51 0.05 4.44 36.51 11.21 -1.69
Variants* 8,133 11.94 0.03 5.37 31.11 5.33 -1.72
Non-COVID 14,848 16.55 -0.07 10.46 36.12 10.31 -1.75
Test-trace 55,893 13.99 -0.07 7.11 37.56 9.83 -1.66
School 13,297 10.02 -0.002 15.90 45.66 10.32 -1.78
Politics 82,746 5.37 -0.05 5.37 31.11 5.33 -1.67
Questioning Institutions** 35,350 2.88 -0.06 5.70 29.20 4.20 -1.71
Race** 8,080 13.48 -0.1 6.5 40.00 13 -1.70
Figure 11. The figure on the left shows overall sentiment distribution in the CDC crisis public. The figure on the right shows sentiment distribution for the School management theme.
Figure 12. The figure on the top left shows sentiment distribution over the Politics theme. The top right figure shows sentiment distribution for the Mitigation theme. The bottom left figure shows sentiment distribution for the Testing theme. Finally, the bottom right figure shows the sentiment distribution for the Non-COVID theme.
Table 16. Themes along with positive and negative top words
Theme Doc2Vec +ve LLR +ve Doc2Vec -ve LLR -ve
Vax booster, pfizer, moderna, eligible, years old, pro-vaccine, vaccines work, #vaccinessavelives, [scientist accounts] vaccine, vaccinated, booster, shot, pfizer, dose, moderna, pregnant,J&J,MRNA vaccinesideeffects, [accounts sharing information about J&J], leaky, experimental, antivxxers, thirdshot, vaccinebooseters, pausing, approving, prove vaccine, flu, unvaccinated, vax, risk, vaers, sever, clot, card, side
Masks mask, wear, face, cloth, protect, spread, help, prevent,respirator, mandate cloth, homemade, fabric, facemask, indoors, breathable, outdoors, layer, [university and scientist accounts] maskitup, againcdc, wtaf, againwith, newyorkgovernor, impeachtrump, masklife, no masks, fucking, unamerican mask, wear, face, droplet, cloth, mandate, ineffective, useless, beard
Variants fully, vaxxed, cloth, homemade, boosted, antivaccers, vaxxup, masking, frustrated,studentathletes vaccinated, mask, fully, delta, wear, indoors, omicron, crowded, honor, guidance mandate, triple boosted, deltavariant, nomasks, againcdc, silly, leaveourkidsalone, unmask, fricking, useless vaccinated, mask, delta, variant, fully, wear, omicron, indoor, contagious, maskless
Schools thrilled, educationsc, skirt, athlete, americadpeds, [govorner accnt],[news accounts], cdph, [teacher union leadership], vaxyourfam school, teacher, student, child, kid, education, safety, classroom, district, open school, ace, child, student, ohio. teacher, institutional, abuse, predatory, sexual trauma, addressing, ace, institutional, teacher union leadership,medically, maltreatment, reopenings, publichealthenemyinchief. homescooled, deathcdc
Not COVID [reporter accounts], cdcfound, hellbentvet, a_million_masks, [cruise company accnts], [politician accnt], cdc_amd cruise, safety, holidays, ship, disenfectant, sanitizer, hiv, thanksgiving, family, organic painadvocate, fightpaindaily, nypdspecialops, gross, [accts pro opioids], [accts of pain advocates], [accts of politicians] pain, opioid, suicide, gun, cuirse, vaping, cancer, isolation, rx dea, quarantine, fentenyl
Table 17. Themes along with positive and negative top words
Theme Doc2Vec +ve LLR +ve Doc2Vec -ve LLR -ve
Politics [politician accnts], cdcfound, [accnts against cruise companies] thank, like, trust, great, good, thanks, god, trump, joke, china idiotinchief, crooked, trumpliesabouteverything, impeached, wrecked, sociopath, notmypresident, inept, deceitful, [news orgs/reporter accts], [politician accts], [self-identified politically active user accts] trump, realdonaltrump, china, cut, eviction, january, fired, obama, moratorium, blame, budget, fuck, pandemic, racism
Test [news organization accts], [public health professional accts], [politician accts], accurate, informed, nurselife, cdcfound, [medical research institution accts] positive, test, coronavirus, testing, covid, test, case, novel, kit, tracing liesaredeadly, trumpcrash, wherearethetests, trumpvirus, absurdity, incompetentinchief, pathological, #fauxnews, deflect, [accnts of self-identified liberal activists], [accnts of liberal leaning politicians]] death, flu, died, test, number, COVID, rate, estimate, testing, season, toll, kit, millions
Mitigation not_that_cdc, [airline accts], antivaccers, vaxyourfam, breathable, [accnts of public health professionals], voteyourfuture, a_million_masks vaccine, mask, wear, help, recommends, safe, protect, hand, safety, effective, vaccintation, please, holiday, prevent basically, triple, antivaxxer, ludicrous, wtaf, cdctravel, dickhead, [airline accnts],vaccineregistration, counterproductive mask, vaccine, wear, sick, risk, unvaccinated people, infected, avoid, sever, contagious, cancer, [acct of airline connected to lowering number of COVID qurantine days], overweight, travel, mandate
Table 18. Themes along with neutral top words
Theme Doc2Vec LLR
Schools curious, daycare, safethansory, scgovpress, economics, [university accnt], openschoolssafely, risky, teleworking, [school district ccnts] school, teacher, student, reopening, kid, child, guidance, classroom, district,inflammatory syndrome
Not COVID unexposed, commits, worstamericanever, uproar, commits, curious, clarifies, encourages, inoculated, recommend cdcgov, via, updated, trumpvirusdeathtoll, trumpisaliar, bidenharrislandslide, guidance, stateofemergency, trumpknew, trumpliedpeopledied
Politics [accnts of journalists & political commentators], wearetheworldforhaiti, funnymemes, [self-identified political advocate accounts] cdcgov, readdonaldtrump, harlem, blackwomen, atlanta, [accts of black celebrities], freecovidtesting, staysafestayhealthy, misinformation
Test tallied, testvirusnow, covidquarantine, unexposed, floridadepartmentofhealth, coronalockdown, [athlete accnts], [sports news accts] coronavidus, case, covid, total, via, total, trimpvirusdeathtoll, florida, updated, trumpisaliar, trumpknew, death, stayhome
Mitigation vax, inoculated, unexposed, contradiction, recomment, safethansorry, vaccineboosters, vaccination drive, [accnts of politicians], [accnts of celebrities] mask, fully, vaccinated, wear, cdc, guidance, booster, done, recommendation, airport
Sample of rich data shared in Tweets
Following is a samples of images from from the dataset, especially focused on PELT change-points as presented in §7.1.
Figure 13. CDC graphic describing the necessary changes in School routines given that the vaccines are less affective against the Delta COVID variant
Figure 14. Message from the CDC showing the importance of vaccinating for the Delta variant.
Figure 15. CDC graph showing that the vaccine is still effective against hospitalization from COVID
Figure 16. Map showing the distribution of variants in different states
Figure 17. CDC graph explaining why new variants are changing masking guidelines even for vaccinated people
Figure 18. CDC graph explaining why the Delta variant is causing changes to masking guidelines even for vaccinated people
Figure 19. This image was shared as a “receipt” showing how the CDC changed the guidelines about who should wear a mask. The image is from CDC guidelines that masks should be used by symptomatic people, caregivers, and healthcare workers.
Figure 20. This image was shared to emphasize how the user thought that recent CDC guidelines to double mask (wear two masks) is an overreaction.
Figure 21. Image shared by a user who was arguing that these guidelines are draconian for thanksgiving and arguing that people need to relax mitigation rules to allow people to have a more “normal” thanksgiving experience.
Figure 22. Image taken from the Miami Herald website introduced to argue that not enough testing is being conducted in Florida. The user was arguing that the smaller number of tests was used to report fewer infections, and thus allow the state to re-open the state even though it might be dangerous.